### v1

# Estimation of AMOC transition probabilities using a machine learning based rare-event algorithm

Abstract
The Atlantic Meridional Overturning Circulation (AMOC) is an important component of the global climate, known to be a tipping element, as it could collapse under global warming. The main objective of this study is to compute the probability that the AMOC collapses within a specified time window, using a rare-event algorithm called Trajectory-Adaptive Multilevel Splitting (TAMS). However, the efficiency and accuracy of TAMS depend on the choice of the score function. Although the definition of the optimal score function, called “committor function” is known, it is impossible in general to compute it a priori. Here, we combine TAMS with a Next-Generation Reservoir Computing technique that estimates the committor function from the data generated by the rare-event algorithm. We test this technique in a stochastic box model of the AMOC for which two types of transition exist, the so-called F(ast)-transitions and S(low)-transitions. Results for the F-transtions compare favorably with those in the literature where a physically-informed score function was used. We show that coupling a rare-event algorithm with machine learning allows for a correct estimation of transition probabilities, transition times, and even transition paths for a wide range of model parameters. We then extend these results to the more difficult problem of S-transitions in the same model. In both cases of F- and S-transitions, we also show how the Next-Generation Reservoir Computing technique can be interpreted to retrieve an analytical estimate of the committor function.
1 Introduction
Tipping elements (Lenton et al., 2008; Armstrong McKay et al., 2022) are subsystems of the climate system that may undergo rapid changes due to global warming. In particular, the Atlantic Meridional Overturning Circulation (AMOC) has been suggested to be in a bistable regime (Stommel, 1961). Since then, the AMOC collapse under freshwater forcing has been studied in various models (Rahmstorf, 1996), and was recently found at the higher end of the model complexity scale (van Westen and Dijkstra, 2023). Observations also suggest that the present-day AMOC is in such a bistable state (Bryden et al., 2011; Garzoli et al., 2013; Weijer et al., 2019; van Westen and Dijkstra, 2023). Considering the high impact of a collapse of the AMOC on the global climate, the probability that it occurs in the coming century should be assessed, even though this transition is a rare event.
Due to its very low likelihood, this event cannot be studied by direct numerical simulations, even in conceptual models. From a theoretical point of view, Large Deviation Theory (Freidlin and Wentzell, 1998) is the natural framework for studying such events and has recently been applied (Soons et al., 2023) to a conceptual AMOC model. For this model (Wood et al., 2019), so-called optimal transition paths (instantons) could be determined, which show the most likely path for the AMOC collapse. In addition, the instantons provide the optimal freshwater forcing needed to effectively collapse the AMOC. However, due partly to the required mathematical assumptions, e.g. low noise, it is not guaranteed that such a method can be applied to fully-coupled climate models.
On the other hand, Trajectory-Adaptive Multilevel Splitting (TAMS) (Lestang et al., 2018) was designed to estimate transition probabilities in fixed time between separate subsets of the phase space. More precisely, given two sets $A$ (e.g., present-day AMOC) and $B$ (e.g., collapsed AMOC), TAMS iteratively drives any trajectory starting in $A$ to reach $B$ within a chosen time, and returns the probability of occurrence of such a transition. This process relies on a score function that discards trajectories unlikely to transition, while cloning and resimulating promising ones without influencing the transition probability. This implies that the score function is key in ensuring the efficiency of TAMS: a bad choice of score function leads to increased variance in the probability estimates, which is also obtained at a (much) larger computational cost. This last point is especially important when applying TAMS to large-dimensional climate models, since their complexity makes simulated data scarce. This class of algorithms has already been successfully applied to climate problems, in particular extreme events, even in a climate model of full complexity (Ragone et al., 2018).
In most cases, a prescribed score function was used in TAMS, although the optimal score function is known to be the committor function (Lestang et al., 2018; Cérou et al., 2019). It is a function of the phase space that associates to any point $\mathbf{x}$ the probability that trajectories starting from $\mathbf{x}$ reach $A$ before $B$. This function is so complex that it is impossible to compute exactly in general, even in simple dynamical systems (Lucente et al., 2022). Fortunately, there are several data-driven methods to estimate the committor function, and these have been compared by Jacques-Dumas et al. (2023): Analogues Markov Chain (Lucente et al., 2022), Feedforward Neural Networks, Next-Generation Reservoir Computing (Gauthier et al., 2021) and Dynamical Galerkin Approximation (Finkel et al., 2021). The Feedforward Neural networks were found to give the best score but the Next-Generation Reservoir Computing (RC) technique could achieve a similar score while requiring less computational time. The RC also presents several other advantages, such as the possibility to train it online and its interpretability.
Here, we combine TAMS with RC to estimate transition probabilities while estimating the committor on-the-fly, using no more data than required by TAMS. We will apply this technique to the same conceptual AMOC model as in Castellana et al. (2019), for which there are two types of transitions: F(ast)-transitions during which the AMOC strength goes to zero due to a noisy freshwater forcing, and S(low)-transitions where the AMOC collapses to a different statistical equilibrium state. As TAMS was already applied to the F-transitions in Castellana et al. (2019) using a prescribed score function, these results serve as a comparison with the TAMS-RC method.
In section 2, we present the conceptual AMOC model, the TAMS algorithm, the committor function, and the RC method with a focus on its combination with TAMS. Then, in section 3 we compare the results of the TAMS-RC method with those in Castellana et al. (2019). In Section 4, we extend the results of Castellana et al. (2019) to compute the probabilities of S-transitions. Finally, in section 5 we show how interpretable information can be extracted from trained neural networks, and in Section 6 we summarize and discuss the results and suggest possible future improvements.
2 Model and methods
2.1 The Cimatoribus-Castellana model
Figure 1: Summarizing picture of the AMOC model, in its version from Castellana et al. (2019). Blue arrows represent the freshwater forcings, red arrows represent the volume transports and cyan arrows stand for wind-driven transports. Solid arrows are always present whatever the AMOC regime, dashed arrows correspond to the present-day AMOC regime and dotted arrows represent the fully collapsed AMOC.
Figure 1: Summarizing picture of the AMOC model, in its version from Castellana et al. (2019). Blue arrows represent the freshwater forcings, red arrows represent the volume transports and cyan arrows stand for wind-driven transports. Solid arrows are always present whatever the AMOC regime, dashed arrows correspond to the present-day AMOC regime and dotted arrows represent the fully collapsed AMOC.
Conceptual models have been very useful for understanding possible multi-stable regimes of the AMOC (Dijkstra, 2023). Here, we use a conceptual AMOC model which was first introduced by Cimatoribus et al. (2014) and slightly modified by Castellana et al. (2019), with the configuration described in Jacques-Dumas et al. (2023). The Atlantic Ocean is represented by five boxes (Fig. 1). There is a northern Atlantic and a southern Atlantic box, respectively, labeled $n$ and $s$. The pycnocline layer is divided into two boxes: a tropical box (labeled $t$) and a southern tropical box (labeled $ts$, located south of $30^{\circ}$). Finally, there is a deep box (labeled $d$) that extends below the other four boxes, from the pycnocline depth to the ocean floor.
In this model, the temperature of boxes n and ts is prescribed and fixed and the state vector consists simply of ($S_{t}$, $S_{t}s$, $S_{n}$, $S_{s}$, $S_{d}$, $D$), the salinity of the boxes and the pycnocline depth $D$. The volume transports between each box are represented by two wind-driven subtropical gyres $r_{s}$ and $r_{n}$, but most importantly by the three quantities: $q_{N}$, $q_{S}$ and $q_{U}$. $q_{N}$ corresponds to the downwelling that takes place in the Northern Atlantic, i.e. the AMOC strength. $q_{S}$ stands for the transport in the South Atlantic and is the difference between the wind-driven Ekman flow ($q_{Ek}$) and the eddy-induced volume transport ($q_{e}$). $q_{U}$ models the Ekman upwelling through the pycnocline.
This model is forced by two separate freshwater fluxes. The first, $E_{s}$, is constant and symmetric from the box $t$ to the boxes $s$ and $n$. The second, $E_{a}$, is asymmetric and is applied from box $s$ to box $n$ and consists of a constant part and a white noise component, multiplied by an amplitude that can be controlled: $E_{a}(t)=\overline{E_{a}}(1+f_{\sigma}\zeta(t))$, where $\zeta(t)$ is a white noise with zero mean and unit variance. The choice of $(\overline{E_{a}},f_{\sigma})$ determines the behavior of the system. The complete equations of the model and the standard configuration of the parameters used are given in Jacques-Dumas et al. (2023) and are not repeated here.
As explained in Castellana et al. (2019), for $\overline{E_{a}}\in[0.06,0.35]$, the AMOC in this model is in a bistable regime. Both its stable fixed points can be characterized in terms of volume transports. The first stable state has $q_{N}>q_{S}>q_{U}>0$; the downwelling in the North Atlantic is much stronger than the upwelling in the South Atlantic and the upwelling through the pycnocline, corresponding to a present-day-like AMOC circulation. The second stable state has $q_{N}=0\ \&\ q_{S}<0$, which means that the downwelling in the North Atlantic is shut down and the circulation in the Southern Hemisphere is reversed; this corresponds to a fully collapsed AMOC. For convenience, we will call these fixed points the “on” state and “off” state of the AMOC, respectively. A transition from the on-state to the off-state is called an S-transition and typically takes about $1,000$ years. Apart from S-transitions, F-transitions also occur in this model, during which the AMOC collapses to a state where $q_{N}=0\ \&\ q_{S}>0$. What happens in this case is that fast freshwater inputs due to noise may shut down the downwelling in the North Atlantic before the deep layer of the ocean can be disturbed, thus maintaining the upwelling in the South Atlantic. Since water accumulates above the pycnocline, this state is only transient, and either the AMOC recovers eventually or completely collapses. The typical timescale for such a temporary shutdown is about $100$ years.
2.2 Trajectory-Adaptive Multilevel Sampling (TAMS)
The general problem to determine the probability of either F- or S-transitions can be rephrased in more general terms as: given two areas of the phase space $A$ and $B$, what is the probability that a trajectory starting in $A$ reaches the zone $B$ before a certain time $T_{\mathrm{max}}$? For example, for an F-transition, $A$ would correspond to a neighborhood of the AMOC-on state, $B$ to $q_{N}=0$ and $T_{\mathrm{max}}=100$ years.
This problem is hard, if not impossible, to treat with direct numerical simulation, especially in the case of rare events when there is a very low probability of observing a transition within a limited time. Splitting methods, on the other hand, are specially designed for this kind of tasks. In particular, the TAMS algorithm (Lestang et al., 2018) is designed to compute the probability of occurrence of rare events at a much lower cost than conventional Monte-Carlo simulations. At every iteration of TAMS, a score function $\Phi$ measures how close each simulated trajectory (among an ensemble) gets to $B$. The ones that are farthest (i.e., lowest score $\Phi$) are discarded and replaced by cloning and branching ”better” trajectories without affecting the transition probability. Thus, at each iteration, all trajectories reach a point that is increasingly close to $B$. When all trajectories have reached $B$, TAMS has converged and outputs an estimate of the transition probability as well as an ensemble of trajectories, each containing a transition.
The steps in the TAMS algorithm are described in Appendix A. Let $q$ be the real transition probability from $A$ to $B$. TAMS should be run $M>1$ times to obtain a distribution of its estimates $\hat{q}$. $\Phi$ is the most important quantity of TAMS, since it controls the deletion and mutation of trajectories and thus drives the trajectories towards $B$. Whatever score function $\Phi$ is used, TAMS is never biased (that is, $\mathbb{E}(\hat{q})=q$), but the variance of the distribution of $\hat{q}$ depends on $\Phi$ (Lestang et al., 2018).
The definition of the optimal score function of the TAMS algorithm, the committor function, is given in Appendix B. For example, for an F-transition in the AMOC model, the set $\mathcal{E}$ in Appendix B is the set $\{\mathbf{x}\in\mathbb{R}^{n}\ |\ q_{N}(\mathbf{x})\leq 0\}$ and the set $\mathcal{F}$ corresponds to $\{(T_{\mathrm{max}},\mathbf{x})\ |\ q_{N}(\mathbf{x})>0\}$. But since TAMS generates trajectories only until $T_{\mathrm{max}}$, the dependence of the committor on the mean passage time $\tau_{\mathcal{F}}$ is automatically implemented and the optimal score function for TAMS now reads (Lestang et al., 2018):
$\Phi(\mathbf{x})=\mathbb{P}(\underset{0\leq t\leq T_{\mathrm{max}}}{\mathrm{min}}q_{N}(\mathbf{X}(t))\leq 0\ |\ \mathbf{X}(0)=\mathbf{x}).$ (1)
Appendix A also gives an expression for the optimal variance $\sigma^{2}_{\mathrm{id}}$ of $\hat{q}$ (when the score function is the committor function).
2.3 Estimating the score function using machine learning for the AMOC model
Our objective is to approach $\Phi$ using the least amount of data possible, that is, by learning this function while running TAMS, using only the data generated by TAMS. In Jacques-Dumas et al. (2023) we studied four data-driven methods that can estimate the committor. We compared their performance in terms of accuracy and computation time and found that the RC method (Gauthier et al., 2021) stood out for complex problems. The RC method (shortly described in Appendix C) is used here with the exact same setup as in Jacques-Dumas et al. (2023).
The combination of TAMS and RC does not pose any major problem: in the code for TAMS, the score function is replaced by the application of the RC to any trajectory where the committor must be computed. What requires more attention are the initialization and training process of the RC. The RC is initialized with $1,000$ trajectories generated for $(\overline{E_{a}}=0.35,f_{\sigma}=0.5)$, where the transition probability is $1$, which makes it easy to obtain transitions, allowing the RC to grasp the structure of the phase space. Due to the lack of data in the low transition probability regime, it ensures that the RC can be initialized at all. During the TAMS run, all discarded trajectories and all trajectories that reach the target set are used to retrain the RC at each iteration. When TAMS finishes, this trained RC is used as the initial score function of TAMS for the parameters $(\overline{E_{a}}=0.35,f_{\sigma}=0.5-\Delta_{f})$ and trained again during the run of TAMS. Setting $\Delta_{f}=0.005$, this process is repeated until $f_{\sigma}=0.005$. For every value of $f_{\sigma}$, the RC trained for $\overline{E_{a}}=0.35$Sv is used as the initial score function for TAMS with $\overline{E_{a}}=0.35\mathrm{Sv}-\Delta_{E_{a}}$ (with $\Delta_{E_{a}}=0.003$) and retrained as explained above. In this way, we cover the entire bistability regime, iteratively adapting the RC to smaller values of $\overline{E_{a}}$. This process is a form of transfer learning: information learned in an easier case (larger noise of forcing leading to a higher transition probability) is transferred to a more difficult case (i.e., lower transition probability) to adapt the neural network faster to this new situation.
Therefore, the RC is constantly retrained during each TAMS run, which means that the quality of the learned score function is highly dependent on the trajectories obtained from TAMS. Since they are subject to noise and randomly cloned, the learned score changes at each TAMS run. To compensate for this effect, we run the combination TAMS-RC $M$ times for each couple of parameters, thus training $M$ independent RC. Each RC is trained using only the trajectories obtained in its own run. This situation is more realistic in the case where we are working with a very large model and we cannot afford to run TAMS many times to optimize the training of the RC.
3 Results: F-transitions
In Castellana et al. (2019), the score function to determine the probability of occurrence of a F-transition of the AMOC was prescribed as:
$\Phi^{F}_{S}(\mathbf{x})=1-\frac{q_{N}(\mathbf{x})}{q^{ON}_{n}},$ (2)
where $\mathbf{x}\in\mathbb{R}^{d}$ is a state in the phase space and the exponent $ON$ refers to the steady on-state of the AMOC model. $\Phi^{F}_{S}(\mathbf{x})$ is equal to $1$ whenever $q_{N}(\mathbf{x})=0$. Here, we define the set $A$ to be a neighborhood of $q^{ON}_{n}$ such that $q_{N}(\mathbf{x})\geq q_{S}(\mathbf{x})$. Given a certain parameter $\rho$, we define $A=\{\mathbf{x}\in\mathbb{R}^{n}\ |\ ||\mathbf{x}-\mathbf{x}^{ON}||\leq\rho\ \&\ q_{N}(\mathbf{x})\geq q_{S}(\mathbf{x})\}$. $\Phi^{F}_{S}(\mathbf{x})$ is set to 00 in this area.
The score function estimated by the RC is called $\Phi_{R}^{F}$. In what follows, the combination of TAMS with the prescribed score function $\Phi^{F}_{S}$ is called TAMS-$S$, and the combination of TAMS with $\Phi_{R}^{F}$ is called TAMS-$R$. We consider $100$ equally spaced values of $\overline{E_{a}}\in[0.06,0.35]$ and $f_{\sigma}\in[0.005,0.5]$. TAMS uses an ensemble of $N=1,000$ trajectories, $n_{c}=10$ of which are discarded at each iteration; $T_{\mathrm{max}}=100$ years. For each couple $(\overline{E_{a}},f_{\sigma})$, TAMS-$S$ and TAMS-$R$ are run both $M=30$ times.
3.1 Consistency indicator $\mathcal{C}$
First, we must make sure that the transition probability estimates obtained with both methods are consistent. We call $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ the mean probability estimates, respectively, obtained with TAMS-$S$ and TAMS-$R$. We call $\sigma_{S}^{F}$ and $\sigma_{R}^{F}$ the corresponding standard deviations. In theory, $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ are equal to the ”true” transition probability (see Sect. 33332.2), but since $N$ and $M$ are finite, they may be biased. It is thus important to ensure that they lie within each other’s confidence interval. We measure the consistency of both distributions by comparing the distance between $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ with the overlap of the $95\%$ confidence intervals attached to both:
$\scriptsize\mathcal{C}=\frac{|\alpha_{S}^{F}-\alpha_{R}^{F}|}{\mathrm{min}\left(\alpha_{S}^{F}+\frac{t\sigma_{S}^{F}}{\sqrt{M}},\alpha_{R}^{F}+\frac{t\sigma_{R}^{F}}{\sqrt{M}}\right)-\mathrm{max}\left(\alpha_{S}^{F}-\frac{t\sigma_{S}^{F}}{\sqrt{M}},\alpha_{R}^{F}-\frac{t\sigma_{R}^{F}}{\sqrt{M}}\right)},$ (3)
where $t$ is the two-sided $95\%$ value from a Student’s distribution. A value of $\mathcal{C}<1$ ensures that both probability estimates are within the confidence interval of each other and thus that no method is biased (Rolland, 2022).
Figure 2: (a) Reproduction of the main result of Castellana et al. (2019). It is a diagram of transition probabilities depending on the parameters $(\overline{E_{a}},f_{\sigma})$ of the model. This diagram was computed with TAMS-$R$.
(b) Consistency score $\mathcal{C}$ for every value of the couple of parameters $(\overline{E_{a}},f_{\sigma})$. The colorbar in this figure is centered on $1$, which means that in blue grid cells, $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ are within each other’s $95\%$ confidence interval. When this is not the case, the grid cell appears in red. Gray cells mean that $\alpha_{S}^{F}$ or $\alpha_{R}^{F}$ was below the $10^{-9}$ cut-off threshold.
The consistency indicator $\mathcal{C}$ is shown in Fig. 2b. $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ are consistent for most values of $(\overline{E_{a}},f_{\sigma})$ since $\mathcal{C}\geq 1$ only in $6.1\%$ of the cases. Fig. 2b can be divided into four regions. In the upper right corner, corresponding to large values of $\overline{E_{a}}$ and $f_{\sigma}$, the transition probabilities are close to $1$ and $\mathcal{C}$ is consistently below $10^{-2}$, which means that $\alpha_{S}^{F}$, $\alpha_{R}^{F}$ and their confidence intervals are almost identical. Then, for lower values of $\overline{E_{a}}$ and $f_{\sigma}$, $\alpha_{R}^{F}$ and $\alpha_{R}^{S}$ are still consistent but the blue show that the confidence intervals do not overlap perfectly anymore. However, $|\alpha_{R}^{F}-\alpha_{R}^{S}|$ remains approximately $10$ times smaller than the overlap of their confidence intervals, so these discrepancies are not significant. Then, the left part of the diagram, whatever the value of $f_{\sigma}$, is dominated by a region where $|\alpha_{S}^{F}-\alpha_{R}^{F}|$ is of the same order of magnitude as the overlap of their confidence interval, although most probability estimates remain consistent. This shows that as $\alpha_{R}^{F}$ decreases, the task of the RC becomes increasingly difficult and the transfer learning process propagates and amplifies biases. Finally, for $f_{\sigma}\leq 0.13$ and the smallest values of $\overline{E_{a}}$ such that $\alpha_{R}^{F}\geq 10^{-9}$, there is a region where inconsistencies dominate. In this regime of extremely low transition probabilities, TAMS-$R$ converges very slowly due to the biases propagated by the very long transfer learning, leading to largely underestimated values of $\alpha_{R}^{F}$. It is striking to note that only $4.5\%$ of the probability estimates $\alpha_{S}^{F}\geq 10^{-6}$ are inconsistent, but this number increases to $34.5\%$ for $\alpha_{S}^{F}<10^{-6}$. However, this also means that the TAMS-$R$ method is very reliable for transition probabilities down to $10^{-6}$, which is lower than often realistically needed.
3.2 Individual variances
$\sigma^{2}_{\mathrm{id}}$ (see equation (10)) is the optimal variance that a given TAMS result must have if $\Phi$ is the committor. If TAMS-$S$ or TAMS-$R$ has provided a certain transition probability, $(\sigma^{2})_{\mathrm{id},\{R,S\}}^{F}$ gives the variance associated with this exact output if it had been obtained with the committor function. So to compare the quality of TAMS-$S$ and TAMS-$R$, we compare the relative difference of their variance to their corresponding ideal variance:
$V_{\mathrm{diff},\{R,S\}}=\frac{|(\sigma^{2})_{\mathrm{id},\{R,S\}}^{F}-(\sigma^{2})_{\{R,S\}}^{F}|}{(\sigma^{2})_{\mathrm{id},\{R,S\}}^{F}}.$ (4)
$V_{\mathrm{diff}}$ is thus a measure of the proximity to the committor: the smaller $V_{\mathrm{diff}}$, the closer $\Phi_{R}^{F}$ or $\Phi_{S}^{S}$ to the committor.
Since the results of TAMS-$R$ are averaged over $30$ independent RCs, to test the variance of this method accurately, we freeze the training of the $30$ RCs for some parameters $(\overline{E_{a}},f_{\sigma})$ and run TAMS-$R$ $30$ times for each of these fixed RCs. We then obtain $30$ variances around the mean probability estimates that are directly comparable to $(\sigma^{2})_{S}^{F}$. We do this for three values of noise amplitude: $f_{\sigma}=0.5,0.25$ and $0.05$. For each of these values of $f_{\sigma}$, we select equally spaced values of $\overline{E_{a}}$ ($1$ out of $10$ in the original $100$ used to compute 2a to explore the entire range of transition probabilities, from $10^{-9}$ to $1$). $V_{\mathrm{diff}}$ is shown in Fig. 3 for the three values of $f_{\sigma}$. On every panel, the blue line corresponds to $V_{\mathrm{diff},S}$ plotted against $\alpha_{S}^{F}$. Each thin black line represents $30$ runs of TAMS-$R_{i}$, where $i\in[1,30]$ represents each independent RC. Thus, each thin black line shows $V_{\mathrm{diff},R_{i}}$ plotted against $\alpha_{R_{i}}^{F}$. The thick black line is the average of the $30$ thin black lines.
Figure 3: Each panel shows, for a fixed $f_{\sigma}$, the relative difference of the variance of $30$ runs of TAMS to the corresponding ideal variance $(\sigma^{2})_{\mathrm{id}_{\{R,S\}}}^{F}$. It is plotted against the corresponding mean probability estimate $\alpha_{\{R,S\}}^{F}$. The blue line corresponds to $V_{\mathrm{diff},S}$ against $\alpha_{S}^{F}$. Each thin black line corresponds to $V_{\mathrm{diff},R_{i,i\in[1,30]}}$ against $\alpha_{R_{i,i\in[1,30]}}^{F}$. The thicker black line represents the average of the thinner black lines.
For all values of $f_{\sigma}$, there is a great variability between the RCs. However, for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$, despite this variability, all RCs are overall consistent with the behavior of $\Phi^{F}_{S}$. This can also be seen with the average of all RCs, which is very consistent with the curve corresponding to $\Phi^{F}_{S}$, both in terms of the mean transition probability (to be read in abscissa) and of the mean distance to the ideal score (to be read in ordinate). However, $\Phi^{F}_{S}$ systematically shows better agreement with the committor function because $V_{\mathrm{diff},S}$ is generally lower.
An interesting case is that of $f_{\sigma}=0.05$ (Fig. 3c). There, all RCs show a significant deviation from the behavior of $\Phi^{F}_{S}$. They also have much greater variability between them than for larger noise values, both in terms of distance from $(\sigma^{2})_{\mathrm{id},R_{i}}^{2}$ and of mean transition probability. The latter also deviates from what is obtained with $\Phi^{F}_{S}$, although both quantities remain consistent on average. This deviation illustrates the limits of transfer learning: the smaller $f_{\sigma}$, the longer the RC has been trained by transferring information learned during easier TAMS runs. When this process takes too long, the RC may be saturated with information and has difficulty adapting to new and more complex situations. Moreover, since all RCs are trained independently, the longer the transfer learning process lasts, the more likely it is to transfer and amplify biases.
3.3 Mean first passage times and trajectories
Finally, it is also important to make sure that the transition paths are not biased. To do this, two measures are considered: the mean first-passage time (MFPT) and the trajectories obtained as output of TAMS. TAMS only simulates trajectories until they reach the target set, so the MFPT can be sampled by looking at the last time step of these trajectories.
Figure 4: a): $1,000$ trajectories have been sampled from $30$ runs of TAMS-$S$. The transition times have been averaged over this ensemble for different values of $(\overline{E_{a}},f_{\sigma})$. This panel shows the mean first-passage times $\mathrm{MFPT}_{S}$ thus obtained, expressed in model years.
b): The MFPT has also been computed for the TAMS runs with the RC, and this panel shows the relative difference $|\mathrm{MFPT}_{S}-\mathrm{MFPT}_{R}|/\mathrm{MFPT}_{S}$, plotted against the mean transition probability $\alpha_{S}^{F}$.
Ensembles of $1,000$ trajectories were saved for several values of $(\overline{E_{a}},f_{\sigma})$, for both TAMS-$S$ and TAMS-$R$, and the MFPTs were averaged over each ensemble. The resulting diagram corresponding to $\Phi^{F}_{S}$ is shown in Fig. 4a. Fig. 4b plots against $\alpha_{S}^{F}$ the relative difference of the MFPTs obtained with $\Phi_{R}^{F}$ and $\Phi^{F}_{S}$. We see in Fig. 4a that the MFPTs lie between $4.2$ model years for the largest value of $\overline{E_{a}}$ and $f_{\sigma}$ to $62$ model years for $\overline{E_{a}}=0.26\mathrm{\ Sv}$ and $f_{\sigma}=0.05$. The distribution of MFPT values shown in Fig. 4a is the one we could expect, with the shortest MFPT corresponding to the largest transition probabilities and long MFPTs associated with very low transition probabilities.
Fig. 4b shows that there is a very good agreement between the MFPT as predicted by $\Phi^{F}_{S}$ and $\Phi_{R}^{F}$. For $\alpha_{S}^{F}$ of the order of $1$, their relative error can reach $12\%$ (in absolute value). But in that case, the transition times are typically shorter than $10$ years, which is too short to consider such an error to be significant; it can simply be due to the stochastic generation of trajectories. For values of $\alpha_{S}^{F}$ down to $10^{-6}$, the relative error in the estimated transition times is lower than $5\%$, therefore both methods give very consistent results. There is only one outlier: there is a $25\%$ relative error on the longest transition time, which corresponds to an overestimation by $15$ years when using $\Phi_{R}^{F}$. In general, the use of $\Phi_{R}^{F}$ systematically overestimates the transition times (by $3$ to $10\%$) for $\alpha_{S}^{F}<10^{-6}$.
Figure 5: Ensemble of $1000$ trajectories obtained with TAMS-$S$ (in blue) and TAMS-$R$ (in red), for $(\overline{E_{a}},f_{\sigma})=(0.288\mathrm{\ Sv},0.05)$ (corresponding to $\alpha_{S}^{F}=4\times 10^{-4}$). We present here the average of the trajectories that reached $B$ between model years $90$ and $100$. The shaded areas represent the difference between the $5^{\mathrm{th}}$ and $95^{\mathrm{th}}$ percentiles in each case.
The last important quality criterion of a TAMS experiment is the trajectories themselves. To do this, we compare the ensembles obtained for $(\overline{E_{a}},f_{\sigma})=(0.288\mathrm{\ Sv},0.05)$ when sampling $1,000$ trajectories from $30$ runs of TAMS-$S$ and $30$ runs of TAMS-$R$. For $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$ (with similar values of $\overline{E_{a}}$), the results are similar as discussed below (not shown). In this case, $\alpha_{S}^{F}=4\times 10^{-4}$. We have chosen these system parameters as a kind of ”worst-case scenario” because the TAMS problem is particularly difficult and the RC is more likely to be imprecise due to the repeated transfer learning processes. Moreover, due to the large variance in the MFPT of the trajectories in this ensemble, we only consider here the trajectories whose transition time is comprised between $90$ and $100$ model years. The average and the interval between the $5^{\mathrm{th}}$ and $95^{\mathrm{th}}$ percentiles of each variable are shown in Fig. 5. The average agreement between the trajectories generated using the score $\Phi_{S}^{F}$ and $\Phi_{R}^{F}$ is very good. For all variables in the model, their averages and the distribution of $90\%$ of their values (throughout the ensemble) closely overlap. These transitions take $90$ to $100$ model years to occur, but we can see on the trajectories that the system stays around its steady on-state for the first $60$ model years.
The physics of the F-transition can also be understood from these trajectories. The transition can only begin when the random freshwater forcing $E_{s}+E_{a}$ is able to overcome the salt advection through the northern subtropical gyre $r_{n}$. Once a random variation of the asymmetric freshwater forcing $E_{a}$ results in a net freshening of the Northern Atlantic (decrease in $S_{n}$), the strength of the downward transport $q_{N}\propto(S_{n}-S_{ts})D^{2}$ also decreases. The salt-advection feedback and the fixed volume of the northern box result in an accumulation of saline water in the tropical box. This accumulation also triggers an increase in $D$, thus an increase of $V_{ts}$ and a freshening of $S_{ts}$. However, this freshening is countered by the increasing effect of the southern subtropical gyre and of the wind-driven Ekman flow (due to the salinification of the tropics). As a result, $S_{ts}$ freshens slower than $S_{n}$, weakening even more $q_{N}$. The increase in $D$ and the effect of the northern tropical gyre are not sufficient to counteract the freshening of the Northern Atlantic and the downwelling stops when $S_{n}$ drops below a critical threshold. Figure 5 shows that during this process, only $S_{n}$, $S_{ts}$ and $D$ vary significantly, validating in hindsight our restriction to these variables for $\Phi_{R}^{F}$.
4 Results: S-transitions
The S-transition case corresponds to a more complex problem for TAMS because the time scale of such a transition is much longer (about $1,000$ model years) and no score function can be derived from physical intuition. Castellana et al. (2019) also mentions this case but did not study it for a wide range of model parameters. For the S-transitions, we study $10$ values of $\overline{E_{a}}\in[0.09,0.345]$ and $f_{\sigma}\in[0.05,0.5]$, spanning the bistability regime. Here, as the equilibria are known, the prescribed score function (Baars et al., 2021) is:
$\Phi^{S}_{S}(\mathbf{x})=\eta-\eta$ $\exp\left(-4\frac{||\mathbf{x}-\mathbf{x}^{ON}||^{2}}{||\mathbf{x}^{OFF}-\mathbf{x}^{ON}||^{2}}\right)$ (5)
$+(1-\eta)$ $\exp\left(-4\frac{||\mathbf{x}-\mathbf{x}^{OFF}||^{2}}{||\mathbf{x}^{OFF}-\mathbf{x}^{ON}||^{2}}\right)$
where |||||$\cdot$||| | ⋅ | | stands for the Euclidean 2-norm and $\eta=||\mathbf{x}^{U}-\mathbf{x}^{ON}||/||\mathbf{x}^{OFF}-\mathbf{x}^{ON}||$ corresponds to the distance between the on-state $\mathbf{x}^{ON}$ and the unstable steady state $\mathbf{x}^{U}$, normalised by the distance between $\mathbf{x}^{ON}$ and the steady off state $\mathbf{x}^{OFF}$. This score function makes sense from a mathematical point of view (the transition between two steady states should go through a saddle-node) but does not allow any physical interpretation. Moreover, it may be strongly biased when $f_{\sigma}$ is large or when the system has multiple unstable steady states. In this case, it is particularly interesting to estimate the committor function via the RC method. We keep the same parameters for TAMS, except for $T_{\mathrm{max}}=1,000$ years, and still call TAMS-$S$ the combination of TAMS with $\Phi_{S}^{S}$ and TAMS-$R$ the combination of TAMS with $\Phi_{R}^{S}$.
TAMS-$R$ is still initialized using an ensemble of $1,000$ pre-generated trajectories with $(\overline{E_{a}}=0.345\mathrm{Sv},f_{\sigma}=0.5)$. But in this case, $\alpha_{S}^{S}=0.078$ so to prevent any bias, we use $500$ trajectories that don’t contain a transition (obtained by direct numerical simulation) $500$ trajectories containing a transition (obtained with TAMS-$S$).
Figure 6: a) Mean probability to transition from $\mathbf{x}^{ON}$ to $\mathbf{x}^{OFF}$ within $1000$ years, over the whole range of values of $(\overline{E_{a}},f_{\sigma})$. This probability was averaged over $30$ runs of TAMS-$S$.
b) Consistency score $\mathcal{C}$ for every value of the couple of parameters $(\overline{E_{a}},f_{\sigma})$, calculated for the transition from $\mathbf{x}^{ON}$ to $\mathbf{x}^{OFF}$. The colorbar in this figure is centered on $1$, which means that, in blue grid cells, $\alpha_{S}^{S}$ and $\alpha_{R}^{S}$ are within each other’s $95\%$ confidence interval. When this is not the case, the grid cell appears in red. Gray cells mean that $\alpha_{S}^{S}$ or $\alpha_{R}^{S}$ is below the $10^{-9}$ cut-off threshold.
Figure 6a shows the diagram of the mean transition probabilities from the AMOC on state to the AMOC off state, computed with $\Phi_{S}^{S}$ and used as reference. Fig. 6b presents the consistency score $\mathcal{C}$ between TAMS-$S$ and TAMS-$R$. Figure 6a shows that, as expected, the smaller the freshwater forcing $\overline{E_{a}}$, the smaller the transition probabilities, as for the F-transitions. However, now a large noise amplitude $f_{\sigma}$ tends to prevent the transition to $\mathbf{x}^{OFF}$. This can be understood by considering the basin of attraction of $\mathbf{x}^{ON}$. Since this basin is very narrow, it only requires a small noise to exit it and be attracted to $\mathbf{x}^{OFF}$. In this setting, a small noise already allows the system to switch basins of attraction and does not disturb it from its natural deterministic drift toward $\mathbf{x}^{OFF}$. On the other hand, a large noise strongly disturbs the system and may force it in directions that do not favor a transition to $\mathbf{x}^{OFF}$, thus lowering the probability of reaching it within $1,000$ years.
The first element to notice in Fig. 6b is that it contains fewer colored grid cells than in 6a. This is due to the fact that the RC was unable to estimate a nonzero probability of collapse when $\alpha_{S}^{S}<10^{-6}$. This failure of the RC for very low transition probabilities may be explained by a less efficient transfer learning process than in the case of the transient collapse. Indeed, the gap between successive sampled parameters $(\overline{E_{a}},f_{\sigma})$ is $10$ times larger than in the previous case, making it more difficult for the RC to adapt to new dynamics. However, this also means that even under these poor learning conditions, the RC can still estimate transition probabilities down to $10^{-6}$ for a wide range of parameters $(\overline{E_{a}},f_{\sigma})$. Namely, the predicted probability $\alpha_{R}^{S}$ is consistent with $\alpha_{S}^{S}$ (so $\mathcal{C}<1$, shown in blue) for $83\%$ of the parameter couples. In all cases where this is not verified (except for one), the inconsistently estimated transition probabilities are the lowest for a given value of $f_{\sigma}$.
Figure 7: (a) Ratio of the number of timesteps computed by TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$S$ require fewer timesteps, blue values mean the opposite.
(b) Ratio of the number of iterations of TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$S$ requires fewer iterations, blue values mean the opposite.
(c) Ratio of the variance in probability estimates obtained with TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$S$ produces a larger variance, blue values mean the opposite. In all panels, the gray color indicates that $\alpha_{S}^{S}$ or $\alpha_{R}^{S}$ is below the $10^{-9}$ cut-off threshold.
Figure 7a compares how many time steps are computed by both methods. Once again, for high probabilities, both methods are, as expected, equivalent because TAMS converges very fast. When they are not equivalent, TAMS-$R$ almost always leads to the computation of more time steps, which is shown in blue in Fig. 7a. The difference is even quite large: in $50\%$ of the cases, TAMS-$R$ leads to the computation of more than $40\%$ more time steps compared to TAMS-$S$. However, if we only consider low probabilities $\alpha_{S}^{S}<10^{-3}$, TAMS-$R$ computes more than $40\%$ more time steps compared to TAMS-$S$ for $40\%$ of the parameter values, and this difference in the calculated time steps is less than $30\%$ of the time steps of TAMS-$S$ for $30\%$ of the parameter values. Therefore, the number of additional time steps computed by TAMS-$R$ compared to TAMS-$S$ stabilizes for lower probabilities.
However, this difference in the number of calculated time steps hides the fact that TAMS-$R$ converges faster than TAMS-$S$. The number of TAMS iterations for both methods is presented in Fig. 7b, and for $90\%$ of the parameters, the number of iterations of TAMS-$R$ is smaller than that of TAMS-$S$. Even when it is larger, the difference is always less than $3\%$ of the number of iterations of TAMS-$R$ (except in two cases where this difference reaches up to $10\%$ of the number of iterations of TAMS-$R$, for probabilities lower than $10^{-4}$). For half of the parameter values, this difference corresponds to more than $15$ more iterations performed by TAMS-$S$ but can reach more than $100$ more iterations in $20\%$ of the parameter values. That TAMS-$R$ performs fewer iterations to converge is consistent with the fact that it computes more time steps: if TAMS converges earlier, it means that it was easier for the trajectories to reach a transition, which implies that they must have been systematically branched at an earlier time. More importantly, a smaller number of iterations also suggests a better score function, since detection of promising and wrong trajectories occurs earlier within the fixed time interval $[0,T_{\mathrm{max}}]$.
Figure 7c represents the variance around the probability estimate for both methods. As is already the case with the other panels, both methods are similar for large transition probabilities: the difference in the values of $(\sigma^{2})^{S}$ is less than $5\%$ of $(\sigma^{2})_{S}^{S}$ only for transition probabilities larger than $0.5$. $(\sigma^{2})_{R}^{S}$ is only smaller than $(\sigma^{2})_{S}^{S}$ for transition probabilities $p\geq 0.08$. Otherwise, TAMS-$S$ systematically yields a smaller variance, with a difference that can be quite large: $\alpha_{R}^{S}$ is more than $50\%$ larger than $\alpha_{S}^{S}$ for half of the parameter values. As is also the case in the other panels of Fig. 7, the smaller the transition probability, the larger this discrepancy and, here, the larger $(\sigma^{2})_{R}^{S}$ compared to $(\sigma^{2})_{S}^{S}$. As we move towards the left of the diagram, the difficulty of the transfer learning does not allow the RC to adapt to new dynamics, and its estimates are increasingly scattered. For even lower probabilities, TAMS-$R$ cannot even converge in many of the $30$ runs but gives a correct probability estimate in others, leading to a very large variance. This also explains why, to the very left of the diagram, there are some cases where $(\sigma^{2})_{R}^{S}<(\sigma^{2})_{S}^{S}$: in this case, there are so many TAMS runs that did not converge and returned a probability of 00 that the prediction of TAMS-$R$ becomes quite consistent. Unlike the variances presented in Fig. 3, the variances shown in Fig. 7c are computed over $30$ runs of TAMS-$R$, each using an independently trained RC.
5 Interpretability of the RC
An important advantage of using RCs is that they can be interpreted. Each column of $\mathbf{W}_{\mathrm{out}}$ (see Appendix C) corresponds to a combination of powers of $S_{n},S_{ts},D$ and $t$. We can thus read in this matrix the terms that contribute the most to the estimated committor.
Figure 8: Each panel shows, for a fixed $f_{\sigma}$, the terms with largest magnitude in the matrix $W_{\mathrm{out}}$ of the RCs. From the $30$ RCs trained in section 3.2, we extract the $10$ terms of $W_{\mathrm{out}}$ with largest magnitude. Their magnitude is then normalized to that of the largest term. We then only keep the terms that are most significant for more than $20$ out of $30$ RCs. Each curve correspond to a value of the parameter $\overline{E_{a}}$ and they show the average magnitude of each term across the $30$ RCs and the errorbars show the corresponding standard deviations.
First, we focus on $\Phi_{R}^{F}$. In the case of F-transitions, we have run TAMS $30$ times for $30$ different RCs at $3$ values of $f_{\sigma}$ (see Sect. 3.2). For each of the RCs, we select the $10$ values of $W_{\mathrm{out}}$ that have the largest amplitude (in absolute value), and we only keep the terms that are common in more than $20$ RCs. Their average magnitude (across the $30$ RCs) is normalized so that the largest is equal to $1$ in absolute value. Figure 8 presents these magnitudes with their standard deviation across the $30$ RCs. For each value of $f_{\sigma}$, if the most significant terms are similar for several values of $\overline{E_{a}}$, they are plotted together on the corresponding panel.
From Fig. 8, we see that for all values of $f_{\sigma}$, the most important terms of all RCs are very consistent: few error bars on the magnitude of every term are visible. This means that all RCs can extract a meaningful structure from the data. Furthermore, for each $f_{\sigma}$, the structure of the matrix $W_{\mathrm{out}}$ also converges as the transition probabilities become very small, as shown by the overlap of the curves obtained for several values of $\overline{E_{a}}$.
These curves also allow for an analytical reconstruction of a suitable score function. For $f_{\sigma}=0.25$ and $f_{\sigma}=0.5$, $3$ interesting terms are highlighted: $S_{n}S_{ts}D^{2}$ (with a magnitude of $-1$) and $S_{n}^{2}D^{2}$ and $S_{ts}^{2}D^{2}$ (with a magnitude of about $0.5$). This naturally leads to a score function $\Phi_{R}^{F}\propto(S_{n}-S_{ts})^{2}D^{2}/2$. In both cases, we can also add the terms $(S_{ts}-S_{n})S_{ts}t^{2}/2$ and $(S_{n}^{2}+S_{ts}^{2}/2-S_{n}S_{ts})Dt/2$. These terms are not exactly contained in the coefficient of $W_{\mathrm{out}}$ because the coefficients of this matrix are not exactly $1$, $0.5$, or $0.25$. These analytical expressions give, nonetheless, a first estimate of an analytical formula for $\Phi_{R}^{F}$, for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$:
$\Phi_{R}^{F}(S_{n},S_{ts},D,t)\propto$ $a_{1}\frac{D^{2}}{2}(S_{n}-S_{ts})^{2}$ (6)
-- $a_{2}\frac{t^{2}}{2}S_{ts}(S_{n}-S_{ts})$
$+$ $a_{3}\frac{Dt}{2}\left(S_{n}^{2}+\frac{S_{ts}^{2}}{2}-S_{n}S_{ts}\right)$
for certain ${\cal O}(1)$ coefficients $a_{i},\ i=1,2,3$, which have appropriate dimensions to make the expression consistent. This expression does not have an ”equal” sign because it is rather a proportionality relation, due to the normalization of $W_{\mathrm{out}}$. $\Phi_{R}^{F}$ can be compared to the structure of $\Phi_{S}^{F}$, driven by the term $(S_{n}-S_{ts})D^{2}$. Here, this precise term cannot be retrieved because the nonlinear part of the feature vector contains only monomials of degree $4$ (see Appendix C). The original implementation of the RC (Gauthier et al., 2021), which we followed, imposes a fixed choice of degree and we found $4$ to work best (Jacques-Dumas et al., 2023). Although $\Phi_{S}^{F}$ cannot be exactly reproduced, the RCs retrieve a similar dependence on $D^{2}$ and $(S_{n}-S_{ts})$. Most importantly, $\Phi_{R}^{F}$ involves time, which is what we should expect from a score function, since the transition must occur before $T_{\mathrm{max}}$. In particular, the third term of Eq. 6 is very interesting, as it combines $D$, $t$, and a term that looks like $(S_{n}-S_{ts})^{2}$, which would not have been possible if the degree of the monomials in the feature vector had been smaller.
Finally, it is striking to notice that the expressions for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$ are very close. This is due in part to transfer learning, but, considering that the RCs are constantly independently retrained, the conservation of this structure across different values of $\overline{E_{a}}$ and $f_{\sigma}$ shows the robustness of this method. However, the RC is still able to adapt, as shown in the third panel of Fig. 8 by the very different structure of $\Phi_{R}^{F}$ for $f_{\sigma}=0.05$. In this case, two monomials stand out: $S_{n}^{2}D^{2}$ and $-S_{ts}^{2}D^{2}$. This suggests a score function $\Phi_{R}^{F}(S_{n},S_{ts},D,t)\propto(S_{n}-S_{ts})(S_{n}+S_{ts})D^{2}$. Other significant terms in these curves are $-(S_{n}^{2}-S_{ts}^{2})Dt/4$ and $(S_{n}^{2}-S_{ts}^{2}/2)t^{2}/2$, leading to a new analytical expression of the score function obtained by the RCs:
$\Phi_{R}^{F}(S_{n},S_{ts},D,t)\propto$ $b_{1}D^{2}(S_{n}^{2}-S_{ts}^{2})$ (7)
-- $b_{2}\frac{Dt}{4}(S_{n}^{2}-S_{ts}^{2})$
-- $b_{3}\frac{t^{2}}{2}\left(S_{n}^{2}-\frac{S_{ts}^{2}}{2}\right)$
for certain ${\cal O}(1)$ coefficients $b_{i},\ i=1,2,3$. The estimated score function for $f_{\sigma}=0.05$ is different from the score obtained for much larger values of the noise amplitude, since a different noise induces different dynamics in the model. Thus, a data-driven score function is expected to depend on the model parameters. If we were able to solve the backward Kolmogorov equation (see Appendix B), the exact committor would depend on $f_{\sigma}$, which would make the structure of the score function evolve with $f_{\sigma}$. However, here such an analytical dependence on $f_{\sigma}$ cannot be retrieved and only its effect can be observed. Finally, despite the different structure of $\Phi_{R}^{F}$ for $f_{\sigma}=0.05$, its dependence on $(S_{n}-S_{ts})$ is conserved, although always multiplied by $(S_{n}+S_{ts})$, stressing more the importance of the salt-advection feedback. The dependence on three terms driven by $D^{2}$, $Dt$ and $t^{2}$ is also conserved.
Such an analysis may be very useful in more complex cases and larger-dimensional models to understand the possible structure of a good score function. We test this point by interpreting the output of $W_{\mathrm{out}}$ when trained to study the transition from the AMOC steady on state to the AMOC steady off-state (studied in Sect. 4).
Figure 9: Each panel shows, for a fixed $f_{\sigma}$, the terms with largest magnitude in the matrix $W_{\mathrm{out}}$ of the RCs. The $10$ dominant terms of $W_{\mathrm{out}}$ are extracted from each of the $30$ RCs. Their magnitude is then normalized to that of the largest term and only terms that are most significant in more than $20$ out of $30$ RCs are retained. Each curve corresponds to a value of the parameter $\overline{E_{a}}$ (so that $\alpha_{R}^{S}$ lies between $0.002$ and $0.1$) and they show the average magnitude of each term across the $30$ RCs and the errorbars show the corresponding standard deviations.
Fig. 9 shows the same type of curves as Fig. 8, obtained from the matrix $W_{\mathrm{out}}$ of $30$ RCs trained on an S-transition of the AMOC, for two values of $f_{\sigma}$: $0.5$ and $0.1$. For each value of $f_{\sigma}$, several curves are presented, corresponding to probability estimates $\alpha_{R}^{S}$ between $0.002$ and $0.1$. For both values of $f_{\sigma}$, all curves are very consistent, showing the ability of the RCs to capture some structure of the phase space and to converge, although they are constantly independently retrained. However, in both panels, the curves corresponding to the lowest transition probabilities have large error bars, showing the limits of transfer learning with a large difference between consecutive values of $\overline{E_{a}}$. In this case, the analytical formulas for the committor that can be extracted from these curves are:
$f_{\sigma}=0.5:\ \Phi_{R}^{S}\propto$ $-c_{1}t^{2}\left(\frac{3S_{n}^{2}}{4}-S_{ts}^{2}+\frac{S_{n}S_{ts}}{2}\right)$ (8)
$-c_{2}\frac{D^{3}}{4}(S_{n}-3S_{ts})$
$+c_{3}\frac{Dt}{2}(S_{n}-S_{ts})\left(D-\frac{S_{n}+S_{ts}}{2}\right)$
$f_{\sigma}=0.1:\ \Phi_{R}^{S}\propto$ $d_{1}D\left(t+\frac{D}{2}\right)((S_{n}+S_{ts})^{2}-S_{n}S_{ts})$
-- $d_{2}\frac{t^{2}}{2}\left(\frac{3S_{n}^{2}}{2}-S_{ts}^{2}\right)$
-- $d_{3}S_{ts}D^{2}t$
for again specific coefficients $c_{i}$ and $d_{i}$. These expressions are difficult to relate to $\Phi_{S}^{S}$ because the latter relied purely on Euclidean distances in phase space rather than on some specific physical variables.
6 Discussion
This is to our knowledge the first study to couple TAMS to a machine learning technique that estimates the committor function on-the-fly. On the one hand, Lucente et al. (2022) had already coupled Adaptive Multilevel Splitting (an algorithm almost similar to TAMS) to another machine learning technique, the Analogue Markov Chain, but the committor had to be estimated beforehand, thus creating an additional computational cost. On the other hand, direct committor function computation using machine learning has already been performed on complex climate data (Miloshevich et al., 2023b), but only with the objective of short-term probabilistic forecast.
We have coupled TAMS to RC to show that the committor function could be learned at no additional cost than what is necessary to compute transition probabilities. We compared our results to a combination of TAMS experiments with a physics-informed score function and showed that the transition probabilities, the mean first-passage times and the transition paths were consistent. In the case of the F-transition, TAMS-$R$ also yields on average a similar variance in the results as TAMS-$S$, except for very low noise. Regarding the S-transition, both methods remain consistent for transition probabilities within $1000$ years down to $10^{-6}$. Furthermore, TAMS-$R$ consistently requires much fewer time steps to converge than TAMS-$S$, suggesting that in this case RCs can learn a more efficient score function than $\Phi_{S}^{S}$. Finally, we showed that analytical formulas can be derived for the learned committor are consistent in the case of the F-transition with the physics-informed score function.
The next step is to study AMOC tipping in models of much higher dimensions, such as Earth Models of Intermediate Complexity (EMICs) or even Global Climate Models (GCMs). Another rare-event algorithm (Moral and Garnier, 2005; Tailleur and Kurchan, 2006; Lecomte and Tailleur, 2007) has already been applied to EMICs (Ragone et al., 2018; Ragone and Bouchet, 2021; Wouters et al., 2023) and to AMOC tipping in particular (Cini et al., 2023). It relies on other selection/deletion rules than a score function, but makes it more difficult to compute transition probabilities. Another interesting data-driven approach of AMOC tipping is the application of the edge-tracking algorithm by Mehling et al. (2024). It aims at finding the critical state through which an AMOC tipping would occur, using a limited number of simulations, and would thus be tractable in higher dimension. This approach provides insight into the dynamics of this system but not on quantities such as the probability of tipping or transition times.
Extending our method to complex climate models raises the question of its scaling with the dimension of the model. For TAMS-$R$, this question is twofold: the feasibility of simulating enough trajectories on the one hand, and the scaling of RC itself, on the other. Regarding the first point, TAMS is not more computationally demanding than any other rare-event algorithm so their recent successes in EMICs (Cini et al., 2023) are promising for TAMS. When it comes to GCMs (van Westen and Dijkstra, 2023), running TAMS does not seem feasible yet.
However, the current implementation of the RC method poses a real scaling problem. Using monomials as a family of basis functions results in a feature vector whose size grows exponentially with the number of variables. We can only consider a very limited number of variables, but it raises in turn the question whether such simplification is sufficient to describe the dynamics of AMOC tipping in a GCM. Miloshevich et al. (2023a) recently showed that Convolutional Neural Networks (CNN) are suitable to learn the committor from maps of climate variables by reducing them into a lower-dimension vector. However, the problem remains that CNNs require a large amount of data to be trained. More generally, this raises the problem of the choice of basis functions for the RC. The same issue arises for another committor estimation method, the Dynamical Galerkin Approximation (Thiede et al., 2019; Finkel et al., 2021; Jacques-Dumas et al., 2023). It is interesting to note that Strahan et al. (2023) recently used a Feedforward Neural Network to solve the Feynman-Kac formula for the committor. This approach only requires short trajectories such as those that can be obtained with TAMS but the issue of the amount of data and the training of the neural network remains.
The interpretability of the RC could also prove useful on the question of dimensionality. For example, we showed in Section 5 that $\mathbf{W}_{\mathrm{out}}$ contains forty basis terms but only about ten play a significant role in the final score function. If these can be detected early enough in the learning process, dropping them would greatly speed up the computation. More generally, neural networks are often considered black boxes and difficult to interpret. Here, it is possible, by estimating analytical formulas, to highlight important variables for the system dynamics and their relationship. It must be used with caution, but can serve as a useful starting point to study the dynamics of the system or to focus observations or experiments. The efficiency of TAMS is considered to rely on sufficient insight into the model’s dynamics (Lestang et al., 2018), but we show here that an efficient TAMS can also be used to obtain such useful insight.
Finally, the coupling between TAMS and RC must be refined, especially the initialization and training process of the latter. In complex models, transition probabilities may only be computed for a single model setup, which makes transfer learning unusable. The RC could then be initialized from a linear score function that makes it easy to generate surrogate data. Since we might not be able to make a better guess of a prescribed score function, this could be a suitable starting point for the learning process. Another important issue is the memory of the RC. The current transfer learning process probably may retain too much information corresponding to very different system parameters, since we did not ”clean” its memory during the computations. This problem does not only arise with transfer learning, and we would also encounter it if the initialization of the RC is very far from the actual committor. A possible solution could be to reinitialize the RC at certain milestones, once a certain number of transitions have been reached, thus using ”better” data.
Acknowledgements.
[…]
  1. 1.
    Initialize $k=0$ and $w_{0}=1$.
  2. 2.
    Generate $N$ trajectories $\left(\mathbf{X}^{(1)},\dots,\mathbf{X}^{(N)}\right)$ starting in set $A$ until $T_{\mathrm{max}}$ or until they reach set $B$. Repeat steps $3-9$ until $N-n_{c}$ trajectories have reached $B$ (or $N$ if $n_{c}=1$).
  3. 3.
    Compute the score function $\Phi(\mathbf{x})$ on every point of each trajectory.
  4. 4.
    For trajectory $i$, the maximum value of the score function is called $Q_{i}$.
  5. 5.
    Let $I_{k}=\{i\in[1,N]\ |\ Q_{i}\in\{Q^{*}_{1},...,Q^{*}_{n_{c}}\}\}$, where $\{Q^{*}_{j}\}_{j\in[1,n_{c}]}$ are the $n_{c}$ lowest values of $Q_{i}$. Since the system is discrete, $I$ may contain more than $n_{c}$ elements.
  6. 6.
    Set $k=k+1$ then $w_{k}=\left(1-\frac{\#I_{k}}{N}\right)w_{k-1}$. For every $i\in I_{k}$, repeat steps $7-9$.
  7. 7.
    Randomly select one of the remaining trajectories, say $\mathbf{X}^{(r)}$ (such that $r\notin I_{k}$). We call $\tau$ the first time so that $\Phi(\mathbf{X}^{(r)}(\tau))=Q_{i}$.
  8. 8.
    Set $\mathbf{X}^{(i)}([1,\tau])=\mathbf{X}^{(r)}([1,\tau])$.
  9. 9.
    Simulate the rest of $\mathbf{X}^{(i)}$ starting from $\mathbf{X}^{(r)}(\tau)$ until time $T_{\mathrm{max}}-\tau$ or until reaching $B$.
We call $N_{B}$ the final number of trajectories that contain a transition, and the total number of iterations is called $K$. The estimated transition probability from $A$ to $B$ before $T_{\mathrm{max}}$ is given by the equation (Rolland and Simonnet, 2015; Rolland, 2022):
$\hat{q}=\frac{N_{B}}{N}w_{K}=\frac{N_{B}}{N}\prod_{k=1}^{K}w_{k}=\frac{N_{B}}{N}\prod_{k=0}^{K-1}\left(1-\frac{\#I_{k}}{N}\right)$ (9)
The optimal variance, obtained when the score function is the committor function (see Appendix B), reads:
$\sigma^{2}_{\mathrm{id}}=\frac{<\hat{q}>^{2}}{N}\left(<K>\frac{n_{c}}{N-n_{c}}+\frac{N-<N_{B}>}{<N_{B}>}\right)$ (10)
where <>$\cdot$>< ⋅ > represents an average over the $M$ runs of TAMS.
[…]
Consider a stochastic dynamical system of dimension $d$ and two subsets $\mathcal{E}$ and $\mathcal{F}$ of its phase space. If the drift of the system is called $a$ and its diffusivity $D$, then the adjoint $\mathcal{L}$ of the Fokker-Planck operator is the infinitesimal generator of the stochastic process and defined as:
$\mathcal{L}=\frac{\partial}{\partial t}(\cdot)-\sum_{i=1}^{n}a_{i}(\mathbf{x})\frac{\partial}{\partial\mathbf{x}_{i}}(\cdot)-\sum_{i,j=1}^{n}D_{ij}(\mathbf{x})\frac{\partial}{\partial\mathbf{x}_{i}\partial\mathbf{x}_{j}}(\cdot),$ (11)
Formally, the committor function, as solution to the backward Kolmogorov equation, solves the Dirichlet problem (Lucente et al., 2022):
$\mathcal{L}\Phi(t,\mathbf{x})$ $=0\ \forall(t,\mathbf{x})\in\mathbb{R}\times(\mathcal{E}\cup\mathcal{F})^{c}$ (12)
$\Phi(t,\mathbf{x})$ $=1\ \forall(t,\mathbf{x})\in\mathbb{R}\times\mathcal{E}$
$\Phi(t,\mathbf{x})$ $=0\ \forall(t,\mathbf{x})\in\mathbb{R}\times\mathcal{F}$
The committor function $\Phi(t,\mathbf{x})$, as defined by the system (12) and its boundary conditions, also corresponds to the probability that a trajectory $\mathbf{X}(t)$ starting on $\mathbf{x}$ reaches the set $\mathcal{E}$ before $\mathcal{F}$. More precisely, for any set $\mathcal{G}\in\mathbb{R}^{n}$, the first passage time in $\mathcal{G}$ is defined as:
$\tau_{\mathcal{G}}=\min\{t\ |\ \mathbf{x}(t)\in\mathcal{G}\}$ (13)
The committor is defined as the probability that the first-passage time in $\mathcal{E}$ is lower than that in $\mathcal{F}$:
$\Phi(t,\mathbf{x})=\mathbb{P}(\tau_{\mathcal{E}}<\tau_{\mathcal{F}}\ |\ \mathbf{X}(0)=\mathbf{x})$ (14)
[…]
Reservoir Computing (Jaeger, 2001; Lukoševičius and Jaeger, 2009) is a kind of neural network that is designed to better represent dynamical systems and temporal processes than classical feedforward neural networks, without using gradient descent for training.
Reservoir Computing aims at transforming a nonlinear problem in the original phase space into a linear problem by linearly embedding input data into a larger-dimensional space. Then a linear regression is performed in this feature space. The main idea is that the matrix representing this space is random and fixed; only the linear regression parameters are learned during training. Although this neural network acts as a universal approximator (Gonon and Ortega, 2020), its use encounters certain problems, such as the randomness of the reservoir: different realizations of a random network can greatly change the performance of the reservoir computer.
To solve this issue, Gauthier et al. (2021) developed another universal approximator (Gonon and Ortega, 2020; Hart et al., 2021) that they call “Next-Generation” Reservoir Computing (here called RC). The difference is that the input data is now embedded into a larger-dimensional space using a family of nonlinear basis functions instead of a random matrix. We provide below only the important steps to build an RC for our problem. Details of the architecture can be found in Gauthier et al. (2021) and in Jacques-Dumas et al. (2023) for the setup and choice of parameters.
At each time step, the RC is given a vector $\mathbf{x}\in\mathbb{R}^{d}$ that contains here only $S_{n}$, $S_{ts}$, $D$ and time $t$. The family of non-linear functions consists of products of power $p$ of these variables. The resulting feature vector $\mathbf{\tilde{x}}\in\mathbb{R}^{M}$, where $M=40$, is much smaller than in classical reservoir computing (where its dimension would be approximately $1,000$), making the training of RC less computationally intensive.
Finally, the output matrix $\mathbf{W}_{\mathrm{out}}\in\mathbb{R}^{2\times M}$ maps $\mathbf{\tilde{x}}$ to the target output $\mathbf{y}\in\mathbb{R}^{2}$. Since we do not have access to any values of the committor, these cannot be used as target output for the linear regression. Instead, the only accessible information is whether $\mathbf{x}$ first leads the trajectory to $A$ or to $B$ or if $T_{\mathrm{max}}$ is reached before any of these sets. So the target output is $\mathbf{y}=(y_{1},y_{2})$ where $y_{1}=1$ if $A$ or $T_{\mathrm{max}}$ is reached first and 00 otherwise ; $y_{2}=1$ if $B$ is reached first and 00 otherwise. A Softmax operation then transforms these predictions finally into probabilities, namely: (”probability to reach $A$ or $T_{\mathrm{max}}$ first”, ”probability to reach $B$ first”). The latter corresponds to the committor function.
Because it is a simple linear regression that transforms $\tilde{\mathbf{x}}$ into $\mathbf{y}$, an analytical expression of $\mathbf{W}_{\mathrm{out}}$ can be derived (Lukoševičius and Jaeger, 2009; Gauthier et al., 2021), only depending on the feature vectors $\mathbf{\tilde{X}}$ and associated labels $\mathbf{Y}$:
$\mathbf{W}_{\mathrm{out}}=\mathbf{Y}\mathbf{\tilde{X}}^{\mathrm{T}}(\mathbf{\tilde{X}}\mathbf{\tilde{X}}^{\mathrm{T}}+\gamma\mathbf{I})^{-1}$ (15)
where $\mathbf{I}$ is the identity matrix and $\gamma$ is the Tikhonov regularization parameter. In our setup, $\mathbf{\tilde{X}}$ depends only on $\gamma=10^{-4}$ and on $p=4$.
Finally, an important advantage of using RC is their ability to be trained online. Here, we would like to use as little data as possible to train the RC, so we use only the data provided by TAMS and retrain the RC at every iteration of the algorithm. Let $\mathbf{\tilde{X}}_{k}$ and $\mathbf{Y}_{k}$ be all the feature vectors and all the labels we used up to iteration $k$. If new data become available, we can add the newly computed feature vectors and labels (say $\mathbf{\tilde{x}}$ and $\mathbf{y}$) as new columns of $\mathbf{\tilde{X}}_{k}$ and $\mathbf{Y}_{k}$. The matrices $\mathbf{Y}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}$ and $\mathbf{\tilde{X}}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}$ can then be easily updated:
$\mathbf{Y}_{k+1}\mathbf{\tilde{X}}_{k+1}^{\mathrm{T}}$ $=\mathbf{Y}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}+\mathbf{y}\mathbf{\tilde{x}}^{\mathrm{T}}$ (16)
$\mathbf{\tilde{X}}_{k+1}\mathbf{\tilde{X}}_{k+1}^{\mathrm{T}}$ $=\mathbf{\tilde{X}}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}+\mathbf{x}\mathbf{\tilde{x}}^{\mathrm{T}}$

### v3

@@
 We then extend these results to the more difficult problem of S-transitions in the same model.
-In both cases of F- and S-transitions, we also show how the Next-Generation Reservoir Computing technique can be interpreted to retrieve an analytical estimate of the committor function.
+In both cases of F-transitions and S-transitions, we also show how the Next-Generation Reservoir Computing technique can be interpreted to retrieve an analytical estimate of the committor function.
 1 Introduction
@@
 In particular, the Atlantic Meridional Overturning Circulation (AMOC) has been suggested to be in a bistable regime (Stommel, 1961).
-Since then, the AMOC collapse under freshwater forcing has been studied in various models (Rahmstorf, 1996), and was recently found at the higher end of the model complexity scale (van Westen and Dijkstra, 2023).
-Observations also suggest that the present-day AMOC is in such a bistable state (Bryden et al., 2011; Garzoli et al., 2013; Weijer et al., 2019; van Westen and Dijkstra, 2023).
-Considering the high impact of a collapse of the AMOC on the global climate, the probability that it occurs in the coming century should be assessed, even though this transition is a rare event.
-Due to its very low likelihood, this event cannot be studied by direct numerical simulations, even in conceptual models.
-From a theoretical point of view, Large Deviation Theory (Freidlin and Wentzell, 1998) is the natural framework for studying such events and has recently been applied (Soons et al., 2023) to a conceptual AMOC model.
-For this model (Wood et al., 2019), so-called optimal transition paths (instantons) could be determined, which show the most likely path for the AMOC collapse.
-In addition, the instantons provide the optimal freshwater forcing needed to effectively collapse the AMOC.
+The AMOC collapse under freshwater forcing has since been studied in various models (Rahmstorf, 1996) and was recently found at the higher end of the model complexity scale (van Westen and Dijkstra, 2023).
+Observations also suggest that the present-day AMOC is in a bistable state (Bryden et al., 2011; Garzoli et al., 2013; Weijer et al., 2019; van Westen and Dijkstra, 2023).
+An AMOC collapse would have a global impact, for example a decrease of the surface temperature over the Northern Hemisphere and especially Europe (Jackson et al., 2015; van Westen et al., 2024; Cini et al., 2024) and a change of rainfall patterns over tropical regions (Jackson et al., 2015; Parsons et al., 2014; Chang et al., 2008).
+Although doubts exist about the possibility of this collapse and the scale of its consequences (Roquet and Wunsch, 2022), the potential impact of such an event calls for an assessment of its probability of occurrence before $2100$.
+Due to its very low likelihood, this event cannot be studied by direct numerical simulations.
+From a theoretical point of view, Large Deviation Theory (Freidlin and Wentzell, 1998) is the natural framework to study it and has recently been applied to a conceptual AMOC model (Soons et al., 2023) for which the instanton, or the most likely path to the collapse, has been computed.
+This method also provides the optimal freshwater forcing needed to effectively collapse the AMOC.
 However, due partly to the required mathematical assumptions, e.g.
-low noise, it is not guaranteed that such a method can be applied to fully-coupled climate models.
-On the other hand, Trajectory-Adaptive Multilevel Splitting (TAMS) (Lestang et al., 2018) was designed to estimate transition probabilities in fixed time between separate subsets of the phase space.
-More precisely, given two sets $A$ (e.g., present-day AMOC) and $B$ (e.g., collapsed AMOC), TAMS iteratively drives any trajectory starting in $A$ to reach $B$ within a chosen time, and returns the probability of occurrence of such a transition.
-This process relies on a score function that discards trajectories unlikely to transition, while cloning and resimulating promising ones without influencing the transition probability.
-This implies that the score function is key in ensuring the efficiency of TAMS: a bad choice of score function leads to increased variance in the probability estimates, which is also obtained at a (much) larger computational cost.
-This last point is especially important when applying TAMS to large-dimensional climate models, since their complexity makes simulated data scarce.
-This class of algorithms has already been successfully applied to climate problems, in particular extreme events, even in a climate model of full complexity (Ragone et al., 2018).
-In most cases, a prescribed score function was used in TAMS, although the optimal score function is known to be the committor function (Lestang et al., 2018; Cérou et al., 2019).
-It is a function of the phase space that associates to any point $\mathbf{x}$ the probability that trajectories starting from $\mathbf{x}$ reach $A$ before $B$.
-This function is so complex that it is impossible to compute exactly in general, even in simple dynamical systems (Lucente et al., 2022).
-Fortunately, there are several data-driven methods to estimate the committor function, and these have been compared by Jacques-Dumas et al.
-(2023): Analogues Markov Chain (Lucente et al., 2022), Feedforward Neural Networks, Next-Generation Reservoir Computing (Gauthier et al., 2021) and Dynamical Galerkin Approximation (Finkel et al., 2021).
-The Feedforward Neural networks were found to give the best score but the Next-Generation Reservoir Computing (RC) technique could achieve a similar score while requiring less computational time.
-The RC also presents several other advantages, such as the possibility to train it online and its interpretability.
-Here, we combine TAMS with RC to estimate transition probabilities while estimating the committor on-the-fly, using no more data than required by TAMS.
-We will apply this technique to the same conceptual AMOC model as in Castellana et al.
-(2019), for which there are two types of transitions: F(ast)-transitions during which the AMOC strength goes to zero due to a noisy freshwater forcing, and S(low)-transitions where the AMOC collapses to a different statistical equilibrium state.
-As TAMS was already applied to the F-transitions in Castellana et al.
-(2019) using a prescribed score function, these results serve as a comparison with the TAMS-RC method.
+low noise, this method is not guaranteed to be applicable to fully-coupled climate models.
+On the other hand, Trajectory-Adaptive Multilevel Splitting (TAMS) (Lestang et al., 2018) was designed to estimate the probability of any transition in fixed time.
+Given two sets $A$ (e.g., present-day AMOC) and $B$ (e.g., collapsed AMOC), TAMS drives any trajectory starting in $A$ towards $B$ within a chosen time.
+This is done with a score function that discards trajectories unlikely to transition, while cloning and resimulating promising ones.
+This function is key in the efficiency of TAMS: a bad score function leads to increased variance in the probability estimates and makes them more costly to compute.
+This is especially important in large-dimensional climate models, whose complexity makes simulated data scarce.
+Similar algorithms have already been successfully applied to climate problems, in particular extreme events, even in a climate model of full complexity (Ragone et al., 2018).
+The optimal score function of TAMS is the committor function (Lestang et al., 2018; Cérou et al., 2019).
+It associates to any state $\mathbf{x}$ in the phase space the probability that trajectories starting from $\mathbf{x}$ reach $A$ before $B$.
+However, computing it exactly is already impossible in simple dynamical systems (Lucente et al., 2022).
+Fortunately, several data-driven methods exist to estimate this function: Analogues Markov Chain (Lucente et al., 2022), Feedforward Neural Networks, Next-Generation Reservoir Computing (Gauthier et al., 2021) and Dynamical Galerkin Approximation (Finkel et al., 2021).
+The Feedforward Neural networks were found to give the best committor estimate (Jacques-Dumas et al., 2023) but the Next-Generation Reservoir Computing (RC) technique achieved a similar performance while requiring less computational time.
+Other advantages of the RC include the possibility to train it online and its interpretability.
+The purpose of Jacques-Dumas et al.
+(2023) was to develop a benchmark for several committor estimation techniques to determine which is best fit for combination with TAMS.
+Here, we combine TAMS with one of them (RC) to compute transition probabilities while estimating the committor on-the-fly, using no more data than required by TAMS.
+We apply this technique to the same conceptual AMOC model as in Castellana et al.
+(2019), for which there are two types of transitions: F(ast)-transitions, noise-driven transitions during which random freshwater forcings drive the AMOC strength to zero, and S(low)-transitions where the AMOC shifts to another statistical equilibrium.
+This model is interesting for our study because these two kinds of collapse develop on different timescales.
+Moreover, Castellana et al.
+(2019) already used TAMS to compute the probability of occurrence of F-transitions.
+These results were obtained using a prescribed score function and are compared to the TAMS-RC method.
 In section 2, we present the conceptual AMOC model, the TAMS algorithm, the committor function, and the RC method with a focus on its combination with TAMS.
@@
 (2019) to compute the probabilities of S-transitions.
-Finally, in section 5 we show how interpretable information can be extracted from trained neural networks, and in Section 6 we summarize and discuss the results and suggest possible future improvements.
+Finally, in section 5 we show how interpretable information can be extracted from the RC, and in Section 6 we summarize and discuss the results and suggest possible future improvements.
 2 Model and methods
@@
 Solid arrows are always present whatever the AMOC regime, dashed arrows correspond to the present-day AMOC regime and dotted arrows represent the fully collapsed AMOC.
-Figure 1: Summarizing picture of the AMOC model, in its version from Castellana et al.
-(2019).
-Blue arrows represent the freshwater forcings, red arrows represent the volume transports and cyan arrows stand for wind-driven transports.
-Solid arrows are always present whatever the AMOC regime, dashed arrows correspond to the present-day AMOC regime and dotted arrows represent the fully collapsed AMOC.
 Conceptual models have been very useful for understanding possible multi-stable regimes of the AMOC (Dijkstra, 2023).
-Here, we use a conceptual AMOC model which was first introduced by Cimatoribus et al.
-(2014) and slightly modified by Castellana et al.
-(2019), with the configuration described in Jacques-Dumas et al.
+Here, we use a conceptual AMOC model introduced by Cimatoribus et al.
+(2014) with the configuration described in Castellana et al.
 (2019).
@@
 1).
-There is a northern Atlantic and a southern Atlantic box, respectively, labeled $n$ and $s$.
+The northern Atlantic and southern Atlantic boxes are respectively labeled $n$ and $s$.
 The pycnocline layer is divided into two boxes: a tropical box (labeled $t$) and a southern tropical box (labeled $ts$, located south of $30^{\circ}$).
-Finally, there is a deep box (labeled $d$) that extends below the other four boxes, from the pycnocline depth to the ocean floor.
-In this model, the temperature of boxes n and ts is prescribed and fixed and the state vector consists simply of ($S_{t}$, $S_{t}s$, $S_{n}$, $S_{s}$, $S_{d}$, $D$), the salinity of the boxes and the pycnocline depth $D$.
-The volume transports between each box are represented by two wind-driven subtropical gyres $r_{s}$ and $r_{n}$, but most importantly by the three quantities: $q_{N}$, $q_{S}$ and $q_{U}$.
-$q_{N}$ corresponds to the downwelling that takes place in the Northern Atlantic, i.e.
+Finally, there is a deep box (labeled $d$) extending below the other boxes, from the pycnocline to the ocean floor.
+The temperature of all boxes is prescribed and fixed so that the state vector $\mathbf{x}=(S_{t},\ S_{ts},\ S_{n},\ S_{s},\ S_{d},\ D)$ consists of the salinity of the boxes and the pycnocline depth $D$.
+The volume transports between each box are represented by two wind-driven subtropical gyres $r_{s}$ and $r_{n}$ and three quantities: $q_{N}$, $q_{S}$ and $q_{U}$.
+The quantity $q_{N}$ corresponds to the downwelling taking place in the Northern Atlantic, i.e.
 the AMOC strength.
-$q_{S}$ stands for the transport in the South Atlantic and is the difference between the wind-driven Ekman flow ($q_{Ek}$) and the eddy-induced volume transport ($q_{e}$).
-$q_{U}$ models the Ekman upwelling through the pycnocline.
-This model is forced by two separate freshwater fluxes.
+The variable $q_{S}$ stands for the transport in the Southern Atlantic and is the difference between the wind-driven Ekman flow ($q_{Ek}$) and the eddy-induced volume transport ($q_{e}$).
+The quantity $q_{U}$ represents the Ekman upwelling through the pycnocline.
+This model is forced by two freshwater fluxes.
 The first, $E_{s}$, is constant and symmetric from the box $t$ to the boxes $s$ and $n$.
-The second, $E_{a}$, is asymmetric and is applied from box $s$ to box $n$ and consists of a constant part and a white noise component, multiplied by an amplitude that can be controlled: $E_{a}(t)=\overline{E_{a}}(1+f_{\sigma}\zeta(t))$, where $\zeta(t)$ is a white noise with zero mean and unit variance.
-The choice of $(\overline{E_{a}},f_{\sigma})$ determines the behavior of the system.
-The complete equations of the model and the standard configuration of the parameters used are given in Jacques-Dumas et al.
-(2023) and are not repeated here.
-As explained in Castellana et al.
-(2019), for $\overline{E_{a}}\in[0.06,0.35]$, the AMOC in this model is in a bistable regime.
-Both its stable fixed points can be characterized in terms of volume transports.
-The first stable state has $q_{N}>q_{S}>q_{U}>0$; the downwelling in the North Atlantic is much stronger than the upwelling in the South Atlantic and the upwelling through the pycnocline, corresponding to a present-day-like AMOC circulation.
-The second stable state has $q_{N}=0\ \&\ q_{S}<0$, which means that the downwelling in the North Atlantic is shut down and the circulation in the Southern Hemisphere is reversed; this corresponds to a fully collapsed AMOC.
-For convenience, we will call these fixed points the “on” state and “off” state of the AMOC, respectively.
+The second, $E_{a}$, is asymmetric and applied from box $s$ to box $n$.
+It models a physical asymmetry between both boxes that controls the strength of the AMOC while keeping the salinity budget closed.
+The asymmetric freshwater flux reads: $E_{a}(t)=\overline{E_{a}}(1+f_{\sigma}\zeta(t))$ where $\overline{E_{a}}$ and $f_{\sigma}$ are constant and $\zeta(t)$ is a white noise with zero mean and unit variance.
+The choice of $(\overline{E_{a}},f_{\sigma})$ thus determines the (stochastic) behavior of the system.
+The complete equations of the model and the standard configuration of the parameters are given in Jacques-Dumas et al.
+(2023) and not repeated here.
+In this model, the AMOC is in a bistable regime for $\overline{E_{a}}\in[0.06,0.35]\mathrm{\ Sv}$ ($1\mathrm{\ Sv}=10^{6}\mathrm{\ m}^{3}$).
+Both its stable fixed states can be characterized in terms of volume transports.
+The first stable state has $q_{N}>q_{S}>q_{U}>0$; the downwelling in the Northern Atlantic is much stronger than the upwelling in the Southern Atlantic and through the pycnocline, corresponding to a present-day-like AMOC circulation.
+The second stable state has $q_{N}=0$ and $q_{S}<0$, such that the downwelling in the Northern Atlantic is shut down and the circulation in the Southern Hemisphere is reversed; this corresponds to a fully collapsed AMOC (see dotted red arrows on Fig.
+1).
+We call these fixed states the “on” state and “off” state of the AMOC, respectively.
 A transition from the on-state to the off-state is called an S-transition and typically takes about $1,000$ years.
-Apart from S-transitions, F-transitions also occur in this model, during which the AMOC collapses to a state where $q_{N}=0\ \&\ q_{S}>0$.
-What happens in this case is that fast freshwater inputs due to noise may shut down the downwelling in the North Atlantic before the deep layer of the ocean can be disturbed, thus maintaining the upwelling in the South Atlantic.
-Since water accumulates above the pycnocline, this state is only transient, and either the AMOC recovers eventually or completely collapses.
+Apart from S-transitions, F-transitions also occur in this model, during which the AMOC collapses to a state where $q_{N}=0$ and $q_{S}>0$.
+In that case, fast noisy freshwater inputs shut down the downwelling in the Northern Atlantic before the deep layer of the ocean can be disturbed, thus maintaining the upwelling in the Southern Atlantic.
+Since water accumulates above the pycnocline, this state is only transient and the AMOC either recovers or completely collapses (leading to an S-transition).
 The typical timescale for such a temporary shutdown is about $100$ years.
 2.2 Trajectory-Adaptive Multilevel Sampling (TAMS)
-The general problem to determine the probability of either F- or S-transitions can be rephrased in more general terms as: given two areas of the phase space $A$ and $B$, what is the probability that a trajectory starting in $A$ reaches the zone $B$ before a certain time $T_{\mathrm{max}}$?
-For example, for an F-transition, $A$ would correspond to a neighborhood of the AMOC-on state, $B$ to $q_{N}=0$ and $T_{\mathrm{max}}=100$ years.
-This problem is hard, if not impossible, to treat with direct numerical simulation, especially in the case of rare events when there is a very low probability of observing a transition within a limited time.
-Splitting methods, on the other hand, are specially designed for this kind of tasks.
-In particular, the TAMS algorithm (Lestang et al., 2018) is designed to compute the probability of occurrence of rare events at a much lower cost than conventional Monte-Carlo simulations.
+The problem of determining the probability of either F-transitions or S-transitions can be rephrased as: given two areas of the phase space $A$ and $B$, what is the probability that a trajectory starting in $A$ reaches $B$ before a time $T_{\mathrm{max}}$?
+For an F-transition, $A$ corresponds to a neighborhood of the AMOC-on state, $B$ to $q_{N}=0$ and $T_{\mathrm{max}}=100$ years.
+This problem is impossible to treat with direct numerical simulation, especially in the case of rare events due to the low transition probability within a limited time.
+However, splitting methods are designed for this kind of task.
+In particular, the TAMS algorithm (Lestang et al., 2018) computes the probability of occurrence of rare events at a much lower cost than Monte-Carlo simulations.
+This algorithm is detailed in Appendix A.
 At every iteration of TAMS, a score function $\Phi$ measures how close each simulated trajectory (among an ensemble) gets to $B$.
-The ones that are farthest (i.e., lowest score $\Phi$) are discarded and replaced by cloning and branching ”better” trajectories without affecting the transition probability.
-Thus, at each iteration, all trajectories reach a point that is increasingly close to $B$.
-When all trajectories have reached $B$, TAMS has converged and outputs an estimate of the transition probability as well as an ensemble of trajectories, each containing a transition.
-The steps in the TAMS algorithm are described in Appendix A.
-Let $q$ be the real transition probability from $A$ to $B$.
-TAMS should be run $M>1$ times to obtain a distribution of its estimates $\hat{q}$.
-$\Phi$ is the most important quantity of TAMS, since it controls the deletion and mutation of trajectories and thus drives the trajectories towards $B$.
-Whatever score function $\Phi$ is used, TAMS is never biased (that is, $\mathbb{E}(\hat{q})=q$), but the variance of the distribution of $\hat{q}$ depends on $\Phi$ (Lestang et al., 2018).
-The definition of the optimal score function of the TAMS algorithm, the committor function, is given in Appendix B.
-For example, for an F-transition in the AMOC model, the set $\mathcal{E}$ in Appendix B is the set $\{\mathbf{x}\in\mathbb{R}^{n}\ |\ q_{N}(\mathbf{x})\leq 0\}$ and the set $\mathcal{F}$ corresponds to $\{(T_{\mathrm{max}},\mathbf{x})\ |\ q_{N}(\mathbf{x})>0\}$.
-But since TAMS generates trajectories only until $T_{\mathrm{max}}$, the dependence of the committor on the mean passage time $\tau_{\mathcal{F}}$ is automatically implemented and the optimal score function for TAMS now reads (Lestang et al., 2018):
-$\Phi(\mathbf{x})=\mathbb{P}(\underset{0\leq t\leq T_{\mathrm{max}}}{\mathrm{min}}q_{N}(\mathbf{X}(t))\leq 0\ |\ \mathbf{X}(0)=\mathbf{x}).$ (1)
-Appendix A also gives an expression for the optimal variance $\sigma^{2}_{\mathrm{id}}$ of $\hat{q}$ (when the score function is the committor function).
-2.3 Estimating the score function using machine learning for the AMOC model
-Our objective is to approach $\Phi$ using the least amount of data possible, that is, by learning this function while running TAMS, using only the data generated by TAMS.
+The farthest ones (i.e., lowest score $\Phi$) are discarded and replaced by cloning and branching ”better” trajectories.
+Thus, at each iteration, all trajectories get closer to $B$.
+TAMS is said to converge when almost all trajectories have reached $B$ (see Appendix A) and it outputs an estimate of the transition probability as well as an ensemble of trajectories, each containing a transition.
+TAMS thus efficiently generates an ensemble of reactive trajectories while allowing to trace back to the original transition probability of the model.
+Let $\alpha$ be the transition probability from $A$ to $B$ within $T_{\mathrm{max}}$.
+Running TAMS $M>1$ times yields a distribution of its estimates $\hat{\alpha}$.
+Whatever score function $\Phi$ is used, $\mathbb{E}(\hat{\alpha})=\alpha$ in the limit of infinite $M$, but the variance of the distribution of $\hat{\alpha}$ depends on $\Phi$ (Lestang et al., 2018).
+The optimal score function for TAMS, which minimizes this variance, is the committor function (see Lestang et al.
+(2018) and Lucente et al.
+(2022) for its definition and properties).
+Appendix A also gives an expression for the optimal variance $\sigma^{2}_{\mathrm{id}}$ of $\hat{\alpha}$ when the committor is used as score function.
+2.3 Estimating the committor function using machine learning for the AMOC model
+We aim at approaching the committor function while using only the data generated by TAMS.
 In Jacques-Dumas et al.
-(2023) we studied four data-driven methods that can estimate the committor.
-We compared their performance in terms of accuracy and computation time and found that the RC method (Gauthier et al., 2021) stood out for complex problems.
-The RC method (shortly described in Appendix C) is used here with the exact same setup as in Jacques-Dumas et al.
+(2023) we studied four data-driven methods that can perform this task.
+In terms of accuracy and computation time, we found Next-Generation Reservoir Computing (RC) (Gauthier et al., 2021) to stand out.
+This method maps the states of a dynamical system onto a larger-dimensional phase space using a family of nonlinear functions.
+Here, we take monomials of degree $p=4$ that are products of powers of $S_{n},S_{ts},D$ and the time $t$.
+Our implementation of this method is presented in Appendix B and with more details in Jacques-Dumas et al.
 (2023).
-The combination of TAMS and RC does not pose any major problem: in the code for TAMS, the score function is replaced by the application of the RC to any trajectory where the committor must be computed.
-What requires more attention are the initialization and training process of the RC.
-The RC is initialized with $1,000$ trajectories generated for $(\overline{E_{a}}=0.35,f_{\sigma}=0.5)$, where the transition probability is $1$, which makes it easy to obtain transitions, allowing the RC to grasp the structure of the phase space.
-Due to the lack of data in the low transition probability regime, it ensures that the RC can be initialized at all.
-During the TAMS run, all discarded trajectories and all trajectories that reach the target set are used to retrain the RC at each iteration.
-When TAMS finishes, this trained RC is used as the initial score function of TAMS for the parameters $(\overline{E_{a}}=0.35,f_{\sigma}=0.5-\Delta_{f})$ and trained again during the run of TAMS.
-Setting $\Delta_{f}=0.005$, this process is repeated until $f_{\sigma}=0.005$.
-For every value of $f_{\sigma}$, the RC trained for $\overline{E_{a}}=0.35$Sv is used as the initial score function for TAMS with $\overline{E_{a}}=0.35\mathrm{Sv}-\Delta_{E_{a}}$ (with $\Delta_{E_{a}}=0.003$) and retrained as explained above.
-In this way, we cover the entire bistability regime, iteratively adapting the RC to smaller values of $\overline{E_{a}}$.
-This process is a form of transfer learning: information learned in an easier case (larger noise of forcing leading to a higher transition probability) is transferred to a more difficult case (i.e., lower transition probability) to adapt the neural network faster to this new situation.
-Therefore, the RC is constantly retrained during each TAMS run, which means that the quality of the learned score function is highly dependent on the trajectories obtained from TAMS.
-Since they are subject to noise and randomly cloned, the learned score changes at each TAMS run.
-To compensate for this effect, we run the combination TAMS-RC $M$ times for each couple of parameters, thus training $M$ independent RC.
-Each RC is trained using only the trajectories obtained in its own run.
-This situation is more realistic in the case where we are working with a very large model and we cannot afford to run TAMS many times to optimize the training of the RC.
+Combining TAMS and RC poses the problems of initialization and the training process, since enough data is needed to obtain reliable predictions.
+The training of the RC is initialized with $1,000$ trajectories generated for $(\overline{E_{a}}=0.35\mathrm{\ Sv},f_{\sigma}=0.5)$, where the transition probability is $1$, which makes it easy to obtain transitions.
+For each TAMS run with parameters $(\overline{E_{a}}=0.35\mathrm{\ Sv},f_{\sigma})$, all discarded trajectories and all trajectories reaching the target set are used to update the RC at each iteration (see Appendix B).
+When TAMS converges, this trained RC is used as the initial score function of TAMS for parameters $(\overline{E_{a}}=0.35\mathrm{\ Sv},f_{\sigma}-\Delta_{f})$ with $\Delta_{f}=0.005$ and this process is repeated until $f_{\sigma}=0.005$.
+For every value of $f_{\sigma}$, the RC trained for $\overline{E_{a}}=0.35\mathrm{\ Sv}$ is used as the initial score function for TAMS with $\overline{E_{a}}=0.35\mathrm{\ Sv}-\Delta_{E_{a}}$ (with $\Delta_{E_{a}}=0.003\mathrm{\ Sv}$) and retrained as explained above.
+In this way, the RC adapts iteratively to parameters for which transitions are increasingly unlikely.
+This process is a form of transfer learning: information learned in an easier case (i.e., larger noise or forcing leading to a large transition probability) is transferred to a more difficult case (i.e., lower transition probability) to adapt the RC faster to this new situation.
+Each RC is trained using only the trajectories obtained in its own run, making them all independent.
+The transfer learning could also be performed by first looping over $\overline{E_{A}}$, then $f_{\sigma}$.
+We tested this alternative setup and found that the results hardly changed overall.
 3 Results: F-transitions
-In Castellana et al. (2019), the score function to determine the probability of occurrence of a F-transition of the AMOC was prescribed as:
+In Castellana et al. (2019), the score function for TAMS is:
 $\Phi^{F}_{S}(\mathbf{x})=1-\frac{q_{N}(\mathbf{x})}{q^{ON}_{N}},$ (1)
 where $\mathbf{x}\in\mathbb{R}^{d}$ is a state in the phase space and the exponent $ON$ refers to the steady on-state of the AMOC model.
-$\Phi^{F}_{S}(\mathbf{x})$ is equal to $1$ whenever $q_{N}(\mathbf{x})=0$.
-Here, we define the set $A$ to be a neighborhood of $q^{ON}_{n}$ such that $q_{N}(\mathbf{x})\geq q_{S}(\mathbf{x})$.
-Given a certain parameter $\rho$, we define $A=\{\mathbf{x}\in\mathbb{R}^{n}\ |\ ||\mathbf{x}-\mathbf{x}^{ON}||\leq\rho\ \&\ q_{N}(\mathbf{x})\geq q_{S}(\mathbf{x})\}$.
+$\Phi^{F}_{S}(\mathbf{x})=1$ whenever $q_{N}(\mathbf{x})=0$.
+Given a certain parameter $\rho$, we define $A$ to be a neighborhood of $q^{ON}_{N}$ such that $q_{N}(\mathbf{x})\geq q_{S}(\mathbf{x})$: $\ A=\{\mathbf{x}\in\mathbb{R}^{n}\ |\ ||\mathbf{x}-\mathbf{x}^{ON}||\leq\rho\ \&\ q_{N}(\mathbf{x})\geq q_{S}(\mathbf{x})\}$.
 $\Phi^{F}_{S}(\mathbf{x})$ is set to 00 in this area.
 The score function estimated by the RC is called $\Phi_{R}^{F}$.
-In what follows, the combination of TAMS with the prescribed score function $\Phi^{F}_{S}$ is called TAMS-$S$, and the combination of TAMS with $\Phi_{R}^{F}$ is called TAMS-$R$.
-We consider $100$ equally spaced values of $\overline{E_{a}}\in[0.06,0.35]$ and $f_{\sigma}\in[0.005,0.5]$.
+We call TAMS-$S$ the combination of TAMS with $\Phi^{F}_{S}$ and TAMS-$R$ the combination of TAMS with $\Phi_{R}^{F}$.
+We sample $100$ equally spaced values of $\overline{E_{a}}\in[0.06,0.35]\mathrm{\ Sv}$ and $f_{\sigma}\in[0.005,0.5]$.
 TAMS uses an ensemble of $N=1,000$ trajectories, $n_{c}=10$ of which are discarded at each iteration; $T_{\mathrm{max}}=100$ years.
 For each couple $(\overline{E_{a}},f_{\sigma})$, TAMS-$S$ and TAMS-$R$ are run both $M=30$ times.
+TAMS is stopped immediately after $\Phi^{F}_{\{R,S\}}=0$, before the circulation in the Southern Atlantic can be reversed, so that we are certain to sample only F-transitions with this method.
 3.1 Consistency indicator $\mathcal{C}$
-First, we must make sure that the transition probability estimates obtained with both methods are consistent.
+First, the transition probability estimates obtained with both methods must be consistent.
 We call $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ the mean probability estimates, respectively, obtained with TAMS-$S$ and TAMS-$R$.
 We call $\sigma_{S}^{F}$ and $\sigma_{R}^{F}$ the corresponding standard deviations.
-In theory, $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ are equal to the ”true” transition probability (see Sect.
-33332.2), but since $N$ and $M$ are finite, they may be biased.
-It is thus important to ensure that they lie within each other’s confidence interval.
-We measure the consistency of both distributions by comparing the distance between $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ with the overlap of the $95\%$ confidence intervals attached to both:
+We measure the consistency of the distributions of $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ by comparing the difference of their mean values to the overlap of their $95\%$ confidence interval:
 $\scriptsize\mathcal{C}=\frac{|\alpha_{S}^{F}-\alpha_{R}^{F}|}{\mathrm{min}\left(\alpha_{S}^{F}+\frac{t\sigma_{S}^{F}}{\sqrt{M}},\alpha_{R}^{F}+\frac{t\sigma_{R}^{F}}{\sqrt{M}}\right)-\mathrm{max}\left(\alpha_{S}^{F}-\frac{t\sigma_{S}^{F}}{\sqrt{M}},\alpha_{R}^{F}-\frac{t\sigma_{R}^{F}}{\sqrt{M}}\right)},$ (2)
 where $t$ is the two-sided $95\%$ value from a Student’s distribution.
-A value of $\mathcal{C}<1$ ensures that both probability estimates are within the confidence interval of each other and thus that no method is biased (Rolland, 2022).
-Figure 2: (a) Reproduction of the main result of Castellana et al.
-(2019).
-It is a diagram of transition probabilities depending on the parameters $(\overline{E_{a}},f_{\sigma})$ of the model.
-This diagram was computed with TAMS-$R$.
+$\mathcal{C}<1$ ensures that both probability estimates are within the confidence interval of each other so that no method is biased (Rolland, 2022).
+Figure 2: (a) Estimated transition probabilities using TAMS-$R$ depending on the model parameters $(\overline{E_{a}},f_{\sigma})$.
+This reproduces the main result of Castellana et al.
+(2019), who used TAMS-$S$.
 (b) Consistency score $\mathcal{C}$ for every value of the couple of parameters $(\overline{E_{a}},f_{\sigma})$.
@@
 Gray cells mean that $\alpha_{S}^{F}$ or $\alpha_{R}^{F}$ was below the $10^{-9}$ cut-off threshold.
-The consistency indicator $\mathcal{C}$ is shown in Fig.
-2b.
+Fig.
+2b shows the consistency indicator $\mathcal{C}$.
 $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ are consistent for most values of $(\overline{E_{a}},f_{\sigma})$ since $\mathcal{C}\geq 1$ only in $6.1\%$ of the cases.
@@
 2b can be divided into four regions.
-In the upper right corner, corresponding to large values of $\overline{E_{a}}$ and $f_{\sigma}$, the transition probabilities are close to $1$ and $\mathcal{C}$ is consistently below $10^{-2}$, which means that $\alpha_{S}^{F}$, $\alpha_{R}^{F}$ and their confidence intervals are almost identical.
-Then, for lower values of $\overline{E_{a}}$ and $f_{\sigma}$, $\alpha_{R}^{F}$ and $\alpha_{R}^{S}$ are still consistent but the blue show that the confidence intervals do not overlap perfectly anymore.
+In the upper right corner, corresponding to large values of $\overline{E_{a}}$ and $f_{\sigma}$, the transition probabilities are close to $1$ and $\mathcal{C}$ is consistently below $10^{-2}$, so $\alpha_{S}^{F}$, $\alpha_{R}^{F}$ and their confidence intervals are almost identical.
+For lower values of $\overline{E_{a}}$ and $f_{\sigma}$, $\alpha_{R}^{F}$ and $\alpha_{R}^{S}$ are still consistent but the lighter blue shading shows that the confidence intervals do not perfectly overlap anymore.
 However, $|\alpha_{R}^{F}-\alpha_{R}^{S}|$ remains approximately $10$ times smaller than the overlap of their confidence intervals, so these discrepancies are not significant.
-Then, the left part of the diagram, whatever the value of $f_{\sigma}$, is dominated by a region where $|\alpha_{S}^{F}-\alpha_{R}^{F}|$ is of the same order of magnitude as the overlap of their confidence interval, although most probability estimates remain consistent.
-This shows that as $\alpha_{R}^{F}$ decreases, the task of the RC becomes increasingly difficult and the transfer learning process propagates and amplifies biases.
+Then, whatever the value of $f_{\sigma}$, in the region containing the smallest values of $\overline{E_{a}}$ such that $\alpha_{R}^{F}>10^{-9}$, all values of $|\alpha_{S}^{F}-\alpha_{R}^{F}|$ are of the same order of magnitude as the overlap of their confidence interval, although most probability estimates remain consistent.
+This shows that as $\alpha_{R}^{F}$ decreases, the task of the RC becomes increasingly difficult and the transfer learning propagates and amplifies biases.
 Finally, for $f_{\sigma}\leq 0.13$ and the smallest values of $\overline{E_{a}}$ such that $\alpha_{R}^{F}\geq 10^{-9}$, there is a region where inconsistencies dominate.
-In this regime of extremely low transition probabilities, TAMS-$R$ converges very slowly due to the biases propagated by the very long transfer learning, leading to largely underestimated values of $\alpha_{R}^{F}$.
+In this regime of extremely low transition probabilities, TAMS-$R$ converges very slowly due to the biases propagated by the very long transfer learning and largely underestimates $\alpha_{R}^{F}$.
 It is striking to note that only $4.5\%$ of the probability estimates $\alpha_{S}^{F}\geq 10^{-6}$ are inconsistent, but this number increases to $34.5\%$ for $\alpha_{S}^{F}<10^{-6}$.
 However, this also means that the TAMS-$R$ method is very reliable for transition probabilities down to $10^{-6}$, which is lower than often realistically needed.
+Collins et al.
+(2013) and Kriegler et al.
+(2009) estimate that the probability that this transition occurs within $100$ years has an upper bound of a few percent, which is many orders of magnitude above the lowest probability that our method can estimate.
+Moreover, this probability increases with global warming (Kriegler et al., 2009; Armstrong McKay et al., 2022).
+Finally, studies (Hall and Ronald, 2001; Li and Born, 2019; Vettoretti et al., 2022) have shown from paleoclimate simulations that such transition has a non-negligible probability, as it can occur spontaneously several times per $10^{4}$ years.
 3.2 Individual variances
-$\sigma^{2}_{\mathrm{id}}$ (see equation (10)) is the optimal variance that a given TAMS result must have if $\Phi$ is the committor.
-If TAMS-$S$ or TAMS-$R$ has provided a certain transition probability, $(\sigma^{2})_{\mathrm{id},\{R,S\}}^{F}$ gives the variance associated with this exact output if it had been obtained with the committor function.
+$(\sigma^{2})_{\mathrm{id},\{R,S\}}^{F}$ (see equation (9)) gives the optimal variance associated with the transition probability computed with TAMS-$S$ or TAMS-$R$ if it had been obtained using the committor function.
 So to compare the quality of TAMS-$S$ and TAMS-$R$, we compare the relative difference of their variance to their corresponding ideal variance:
 $V_{\mathrm{diff},\{R,S\}}=\frac{|(\sigma^{2})_{\mathrm{id},\{R,S\}}^{F}-(\sigma^{2})_{\{R,S\}}^{F}|}{(\sigma^{2})_{\mathrm{id},\{R,S\}}^{F}}.$ (3)
-$V_{\mathrm{diff}}$ is thus a measure of the proximity to the committor: the smaller $V_{\mathrm{diff}}$, the closer $\Phi_{R}^{F}$ or $\Phi_{S}^{S}$ to the committor.
-Since the results of TAMS-$R$ are averaged over $30$ independent RCs, to test the variance of this method accurately, we freeze the training of the $30$ RCs for some parameters $(\overline{E_{a}},f_{\sigma})$ and run TAMS-$R$ $30$ times for each of these fixed RCs.
-We then obtain $30$ variances around the mean probability estimates that are directly comparable to $(\sigma^{2})_{S}^{F}$.
+$V_{\mathrm{diff}}$ is a measure of the proximity to the committor: the smaller $V_{\mathrm{diff}}$, the closer $\Phi_{R}^{F}$ or $\Phi_{S}^{S}$ to the committor.
+In Fig.
+2a, TAMS-$R$ was run $30$ times using $30$ independent RCs acting as $30$ different score functions.
+The variance over these runs cannot be directly compared to the variance over the $30$ runs of TAMS-$S$, which uses a fixed score function.
+When running TAMS-$R$, the RCs are updated at each TAMS iteration, adapting them to each couple of parameters $(\overline{E_{a}},f_{\sigma})$.
+These trained RCs can then be “frozen” (their parameters are fixed and not retrained anymore) and be used as fixed score function for TAMS-$R$, so that the variance on the TAMS results when using RCs can be compared to the variance of TAMS-$S$.
+TAMS-$R$ is thus run $30$ times again using each of the $30$ independent RCs trained previously ($900$ runs of TAMS in total).
 We do this for three values of noise amplitude: $f_{\sigma}=0.5,0.25$ and $0.05$.
-For each of these values of $f_{\sigma}$, we select equally spaced values of $\overline{E_{a}}$ ($1$ out of $10$ in the original $100$ used to compute 2a to explore the entire range of transition probabilities, from $10^{-9}$ to $1$).
+For each of these values of $f_{\sigma}$, results are saved for one value of $\overline{E_{a}}$ out of $10$ across the whole range $[0.06,0.35]\mathrm{\ Sv}$ (the computation is stopped when $\alpha_{R}^{F}\leq 10^{-9}$).
 $V_{\mathrm{diff}}$ is shown in Fig.
@@
 For all values of $f_{\sigma}$, there is a great variability between the RCs.
-However, for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$, despite this variability, all RCs are overall consistent with the behavior of $\Phi^{F}_{S}$.
-This can also be seen with the average of all RCs, which is very consistent with the curve corresponding to $\Phi^{F}_{S}$, both in terms of the mean transition probability (to be read in abscissa) and of the mean distance to the ideal score (to be read in ordinate).
-However, $\Phi^{F}_{S}$ systematically shows better agreement with the committor function because $V_{\mathrm{diff},S}$ is generally lower.
-An interesting case is that of $f_{\sigma}=0.05$ (Fig.
-3c).
-There, all RCs show a significant deviation from the behavior of $\Phi^{F}_{S}$.
-They also have much greater variability between them than for larger noise values, both in terms of distance from $(\sigma^{2})_{\mathrm{id},R_{i}}^{2}$ and of mean transition probability.
-The latter also deviates from what is obtained with $\Phi^{F}_{S}$, although both quantities remain consistent on average.
-This deviation illustrates the limits of transfer learning: the smaller $f_{\sigma}$, the longer the RC has been trained by transferring information learned during easier TAMS runs.
+However, for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$, all RCs are overall consistent with $\Phi^{F}_{S}$.
+The average of all RCs is also consistent with the curve corresponding to $\Phi^{F}_{S}$, both in terms of the mean transition probability (to be read in abscissa) and of the mean distance to the ideal score (to be read in ordinate).
+However, $\Phi^{F}_{S}$ shows better agreement with the committor function as $V_{\mathrm{diff},S}$ is generally lower.
+For $f_{\sigma}=0.05$ (Fig.
+3c), all RCs show a significant deviation from $\Phi^{F}_{S}$ and have a much greater variability between them in terms of distance to $(\sigma^{2})_{\mathrm{id},R_{i}}^{F}$ than for larger $f_{\sigma}$.
+The lowest values of $\alpha_{S}^{F}$ and $\alpha_{R}^{F}$ also deviate (the left-end of the curves are not at the same point on the horizontal axis), although they remain consistent on average (thick black and blue lines) at about $10^{-8}$.
+This deviation illustrates the limits of transfer learning: the smaller $f_{\sigma}$, the longer the RC was trained by transferring information learned during easier TAMS runs.
 When this process takes too long, the RC may be saturated with information and has difficulty adapting to new and more complex situations.
-Moreover, since all RCs are trained independently, the longer the transfer learning process lasts, the more likely it is to transfer and amplify biases.
+Moreover, since all RCs are independent, the longer the transfer learning process, the more likely it is to amplify biases.
 3.3 Mean first passage times and trajectories
-Finally, it is also important to make sure that the transition paths are not biased.
-To do this, two measures are considered: the mean first-passage time (MFPT) and the trajectories obtained as output of TAMS.
-TAMS only simulates trajectories until they reach the target set, so the MFPT can be sampled by looking at the last time step of these trajectories.
+Finally, the transition paths must not be biased.
+Two measures are considered: the mean first-passage time (MFPT) and the trajectories obtained as output of TAMS.
+TAMS only simulates trajectories until they reach the target set, so the MFPT can be estimated from the last time step of these trajectories.
 Figure 4: a): $1,000$ trajectories have been sampled from $30$ runs of TAMS-$S$.
@@
 4b plots against $\alpha_{S}^{F}$ the relative difference of the MFPTs obtained with $\Phi_{R}^{F}$ and $\Phi^{F}_{S}$.
-We see in Fig.
-4a that the MFPTs lie between $4.2$ model years for the largest value of $\overline{E_{a}}$ and $f_{\sigma}$ to $62$ model years for $\overline{E_{a}}=0.26\mathrm{\ Sv}$ and $f_{\sigma}=0.05$.
-The distribution of MFPT values shown in Fig.
-4a is the one we could expect, with the shortest MFPT corresponding to the largest transition probabilities and long MFPTs associated with very low transition probabilities.
 Fig.
-4b shows that there is a very good agreement between the MFPT as predicted by $\Phi^{F}_{S}$ and $\Phi_{R}^{F}$.
+4a shows that the MFPTs lie between $4.2$ model years for the largest value of $\overline{E_{a}}$ and $f_{\sigma}$ to $62$ model years for $\overline{E_{a}}=0.26\mathrm{\ Sv}$ and $f_{\sigma}=0.05$.
+This distribution of MFPTs is as expected, with the shortest MFPTs corresponding to the largest transition probabilities and long MFPTs associated with low transition probabilities.
+Fig.
+4b shows that the MFPTs predicted by $\Phi^{F}_{S}$ and $\Phi_{R}^{F}$ are in agreement.
 For $\alpha_{S}^{F}$ of the order of $1$, their relative error can reach $12\%$ (in absolute value).
-But in that case, the transition times are typically shorter than $10$ years, which is too short to consider such an error to be significant; it can simply be due to the stochastic generation of trajectories.
-For values of $\alpha_{S}^{F}$ down to $10^{-6}$, the relative error in the estimated transition times is lower than $5\%$, therefore both methods give very consistent results.
-There is only one outlier: there is a $25\%$ relative error on the longest transition time, which corresponds to an overestimation by $15$ years when using $\Phi_{R}^{F}$.
+But in that case, transition times are typically shorter than $10$ years, so such an error can simply be due to the stochastic generation of trajectories.
+For values of $\alpha_{S}^{F}$ down to $10^{-6}$, the relative error in the estimated transition times is lower than $5\%$, therefore both methods are very consistent.
+The only outlier is the $25\%$ relative error on the longest transition time, which corresponds to an overestimation by $15$ years when using $\Phi_{R}^{F}$.
 In general, the use of $\Phi_{R}^{F}$ systematically overestimates the transition times (by $3$ to $10\%$) for $\alpha_{S}^{F}<10^{-6}$.
@@
 The shaded areas represent the difference between the $5^{\mathrm{th}}$ and $95^{\mathrm{th}}$ percentiles in each case.
-The last important quality criterion of a TAMS experiment is the trajectories themselves.
-To do this, we compare the ensembles obtained for $(\overline{E_{a}},f_{\sigma})=(0.288\mathrm{\ Sv},0.05)$ when sampling $1,000$ trajectories from $30$ runs of TAMS-$S$ and $30$ runs of TAMS-$R$.
+Then, we compare the trajectory ensembles obtained for $(\overline{E_{a}},f_{\sigma})=(0.288\mathrm{\ Sv},0.05)$ when sampling $1,000$ trajectories from $30$ runs of TAMS-$S$ and $30$ runs of TAMS-$R$.
 For $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$ (with similar values of $\overline{E_{a}}$), the results are similar as discussed below (not shown).
 In this case, $\alpha_{S}^{F}=4\times 10^{-4}$.
-We have chosen these system parameters as a kind of ”worst-case scenario” because the TAMS problem is particularly difficult and the RC is more likely to be imprecise due to the repeated transfer learning processes.
-Moreover, due to the large variance in the MFPT of the trajectories in this ensemble, we only consider here the trajectories whose transition time is comprised between $90$ and $100$ model years.
+We have chosen these system parameters as a ”worst-case scenario” because the TAMS problem is particularly difficult and the RC is more likely to be imprecise due to the repeated transfer learning processes.
+Moreover, due to the large variance in the MFPTs of the trajectories in this ensemble, we only consider here the trajectories whose transition time lies between $90$ and $100$ model years.
+This corresponds to $124$ trajectories for TAMS-$S$ and $145$ for TAMS-$R$ because, in TAMS, trajectories are left free to collapse at any time during the $100$ simulated years.
+We restricted ourselves to this decade to show typical collapses from trajectories within a full century.
 The average and the interval between the $5^{\mathrm{th}}$ and $95^{\mathrm{th}}$ percentiles of each variable are shown in Fig.
@@
 These transitions take $90$ to $100$ model years to occur, but we can see on the trajectories that the system stays around its steady on-state for the first $60$ model years.
-The physics of the F-transition can also be understood from these trajectories.
-The transition can only begin when the random freshwater forcing $E_{s}+E_{a}$ is able to overcome the salt advection through the northern subtropical gyre $r_{n}$.
-Once a random variation of the asymmetric freshwater forcing $E_{a}$ results in a net freshening of the Northern Atlantic (decrease in $S_{n}$), the strength of the downward transport $q_{N}\propto(S_{n}-S_{ts})D^{2}$ also decreases.
-The salt-advection feedback and the fixed volume of the northern box result in an accumulation of saline water in the tropical box.
-This accumulation also triggers an increase in $D$, thus an increase of $V_{ts}$ and a freshening of $S_{ts}$.
-However, this freshening is countered by the increasing effect of the southern subtropical gyre and of the wind-driven Ekman flow (due to the salinification of the tropics).
-As a result, $S_{ts}$ freshens slower than $S_{n}$, weakening even more $q_{N}$.
-The increase in $D$ and the effect of the northern tropical gyre are not sufficient to counteract the freshening of the Northern Atlantic and the downwelling stops when $S_{n}$ drops below a critical threshold.
-Figure 5 shows that during this process, only $S_{n}$, $S_{ts}$ and $D$ vary significantly, validating in hindsight our restriction to these variables for $\Phi_{R}^{F}$.
 4 Results: S-transitions
@@
 (2019) also mentions this case but did not study it for a wide range of model parameters.
-For the S-transitions, we study $10$ values of $\overline{E_{a}}\in[0.09,0.345]$ and $f_{\sigma}\in[0.05,0.5]$, spanning the bistability regime.
-Here, as the equilibria are known, the prescribed score function (Baars et al., 2021) is:
+For the S-transitions, we consider $10$ equally spaced values of $\overline{E_{a}}\in[0.09,0.345]\mathrm{\ Sv}$ (spanning the bistability regime) and $f_{\sigma}\in[0.05,0.5]$.
+Here, as the equilibria are known, the prescribed score function (Baars et al., 2021; Castellana et al., 2019) is:
 $\Phi^{S}_{S}(\mathbf{x})=\eta-\eta$ $\exp\left(-4\frac{||\mathbf{x}-\mathbf{x}^{ON}||^{2}}{||\mathbf{x}^{OFF}-\mathbf{x}^{ON}||^{2}}\right)$ (4)
 $+(1-\eta)$ $\exp\left(-4\frac{||\mathbf{x}-\mathbf{x}^{OFF}||^{2}}{||\mathbf{x}^{OFF}-\mathbf{x}^{ON}||^{2}}\right)$
-where |||||$\cdot$||| | ⋅ | | stands for the Euclidean 2-norm and $\eta=||\mathbf{x}^{U}-\mathbf{x}^{ON}||/||\mathbf{x}^{OFF}-\mathbf{x}^{ON}||$ corresponds to the distance between the on-state $\mathbf{x}^{ON}$ and the unstable steady state $\mathbf{x}^{U}$, normalised by the distance between $\mathbf{x}^{ON}$ and the steady off state $\mathbf{x}^{OFF}$.
+where |||||$\cdot$||| | ⋅ | | stands for the Euclidean 2-norm and $\eta=||\mathbf{x}^{U}-\mathbf{x}^{ON}||/||\mathbf{x}^{OFF}-\mathbf{x}^{ON}||$ corresponds to the distance between the on-state $\mathbf{x}^{ON}$ and the unstable steady state (saddle-node) $\mathbf{x}^{U}$, normalized by the distance between $\mathbf{x}^{ON}$ and the steady off state $\mathbf{x}^{OFF}$.
 This score function makes sense from a mathematical point of view (the transition between two steady states should go through a saddle-node) but does not allow any physical interpretation.
 Moreover, it may be strongly biased when $f_{\sigma}$ is large or when the system has multiple unstable steady states.
-In this case, it is particularly interesting to estimate the committor function via the RC method.
+In this case, it is particularly interesting to estimate the committor function with the RC.
 We keep the same parameters for TAMS, except for $T_{\mathrm{max}}=1,000$ years, and still call TAMS-$S$ the combination of TAMS with $\Phi_{S}^{S}$ and TAMS-$R$ the combination of TAMS with $\Phi_{R}^{S}$.
-TAMS-$R$ is still initialized using an ensemble of $1,000$ pre-generated trajectories with $(\overline{E_{a}}=0.345\mathrm{Sv},f_{\sigma}=0.5)$.
-But in this case, $\alpha_{S}^{S}=0.078$ so to prevent any bias, we use $500$ trajectories that don’t contain a transition (obtained by direct numerical simulation) $500$ trajectories containing a transition (obtained with TAMS-$S$).
-Figure 6: a) Mean probability to transition from $\mathbf{x}^{ON}$ to $\mathbf{x}^{OFF}$ within $1000$ years, over the whole range of values of $(\overline{E_{a}},f_{\sigma})$.
-This probability was averaged over $30$ runs of TAMS-$S$.
+TAMS-$R$ is initialized using an ensemble of $1,000$ pre-generated trajectories with $(\overline{E_{a}}=0.345\mathrm{\ Sv},f_{\sigma}=0.5)$.
+But in this case, $\alpha_{S}^{S}=0.078$ so to prevent any bias, $500$ trajectories don’t contain a transition (obtained by direct numerical simulation) and $500$ trajectories contain one (obtained with TAMS-$S$).
+Otherwise, due to the rarity of transitions, $1000$ trajectories would contain too few of them and such unbalanced training dataset would hinder performance.
+Figure 6: a) Mean estimated probability to transition from $\mathbf{x}^{ON}$ to $\mathbf{x}^{OFF}$ within $1000$ years, over the whole range of values of $(\overline{E_{a}},f_{\sigma})$.
+This probability was averaged over $30$ runs of TAMS-$R$.
 b) Consistency score $\mathcal{C}$ for every value of the couple of parameters $(\overline{E_{a}},f_{\sigma})$, calculated for the transition from $\mathbf{x}^{ON}$ to $\mathbf{x}^{OFF}$.
@@
 Gray cells mean that $\alpha_{S}^{S}$ or $\alpha_{R}^{S}$ is below the $10^{-9}$ cut-off threshold.
-Figure 6a shows the diagram of the mean transition probabilities from the AMOC on state to the AMOC off state, computed with $\Phi_{S}^{S}$ and used as reference.
+Figure 6a shows the diagram of the mean transition probabilities from the AMOC on state to the AMOC off state, computed with $\Phi_{R}^{S}$.
 Fig.
@@
 Figure 6a shows that, as expected, the smaller the freshwater forcing $\overline{E_{a}}$, the smaller the transition probabilities, as for the F-transitions.
-However, now a large noise amplitude $f_{\sigma}$ tends to prevent the transition to $\mathbf{x}^{OFF}$.
+However, now, a large noise amplitude $f_{\sigma}$ tends to prevent the transition to $\mathbf{x}^{OFF}$.
 This can be understood by considering the basin of attraction of $\mathbf{x}^{ON}$.
@@
 On the other hand, a large noise strongly disturbs the system and may force it in directions that do not favor a transition to $\mathbf{x}^{OFF}$, thus lowering the probability of reaching it within $1,000$ years.
-The first element to notice in Fig.
-6b is that it contains fewer colored grid cells than in 6a.
-This is due to the fact that the RC was unable to estimate a nonzero probability of collapse when $\alpha_{S}^{S}<10^{-6}$.
+The cut-off to compute transition probabilities with TAMS was here as well set to $10^{-9}$ but TAMS-$R$ reaches such a low probability only for $(\overline{E_{a}}=0.15\ \mathrm{Sv},f_{\sigma}=0.25)$.
+In all other cases, it fails to compute transition probabilities lower than $10^{-6}$.
 This failure of the RC for very low transition probabilities may be explained by a less efficient transfer learning process than in the case of the transient collapse.
@@
 However, this also means that even under these poor learning conditions, the RC can still estimate transition probabilities down to $10^{-6}$ for a wide range of parameters $(\overline{E_{a}},f_{\sigma})$.
-Namely, the predicted probability $\alpha_{R}^{S}$ is consistent with $\alpha_{S}^{S}$ (so $\mathcal{C}<1$, shown in blue) for $83\%$ of the parameter couples.
+Fig.
+6b demonstrates that the predicted probability $\alpha_{R}^{S}$ is consistent with $\alpha_{S}^{S}$ (so $\mathcal{C}<1$, shown in blue) for $83\%$ of the parameter couples.
 In all cases where this is not verified (except for one), the inconsistently estimated transition probabilities are the lowest for a given value of $f_{\sigma}$.
-Figure 7: (a) Ratio of the number of timesteps computed by TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$S$ require fewer timesteps, blue values mean the opposite.
-(b) Ratio of the number of iterations of TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$S$ requires fewer iterations, blue values mean the opposite.
-(c) Ratio of the variance in probability estimates obtained with TAMS-$S$ compared to TAMS-$R$.
-Red values mean that TAMS-$S$ produces a larger variance, blue values mean the opposite.
+Figure 7: (a) Ratio of the number of timesteps computed by TAMS-$S$ compared to TAMS-$R$.
+Red values mean that TAMS-$R$ require fewer timesteps, blue values mean the opposite.
 In all panels, the gray color indicates that $\alpha_{S}^{S}$ or $\alpha_{R}^{S}$ is below the $10^{-9}$ cut-off threshold.
+(b) Ratio of the number of iterations of TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$R$ requires fewer iterations, blue values mean the opposite.
+(c) Ratio of the variance in probability estimates obtained with TAMS-$S$ compared to TAMS-$R$. Red values mean that TAMS-$S$ produces a larger variance, blue values mean the opposite.
 Figure 7a compares how many time steps are computed by both methods.
-Once again, for high probabilities, both methods are, as expected, equivalent because TAMS converges very fast.
+Since each iteration of TAMS consists in integrating trajectories, i.e.
+simulating many time steps, we summed them up during the whole TAMS process to obtain an indicator of its computational cost.
+For high probabilities, both methods are, as expected, equivalent because TAMS converges very fast.
 When they are not equivalent, TAMS-$R$ almost always leads to the computation of more time steps, which is shown in blue in Fig.
@@
 The difference is even quite large: in $50\%$ of the cases, TAMS-$R$ leads to the computation of more than $40\%$ more time steps compared to TAMS-$S$.
-However, if we only consider low probabilities $\alpha_{S}^{S}<10^{-3}$, TAMS-$R$ computes more than $40\%$ more time steps compared to TAMS-$S$ for $40\%$ of the parameter values, and this difference in the calculated time steps is less than $30\%$ of the time steps of TAMS-$S$ for $30\%$ of the parameter values.
-Therefore, the number of additional time steps computed by TAMS-$R$ compared to TAMS-$S$ stabilizes for lower probabilities.
-However, this difference in the number of calculated time steps hides the fact that TAMS-$R$ converges faster than TAMS-$S$.
+However, the smaller number of time steps computed by TAMS-$S$ hides the fact that the latter takes more iterations to converge than TAMS-$R$.
 The number of TAMS iterations for both methods is presented in Fig.
@@
 For half of the parameter values, this difference corresponds to more than $15$ more iterations performed by TAMS-$S$ but can reach more than $100$ more iterations in $20\%$ of the parameter values.
-That TAMS-$R$ performs fewer iterations to converge is consistent with the fact that it computes more time steps: if TAMS converges earlier, it means that it was easier for the trajectories to reach a transition, which implies that they must have been systematically branched at an earlier time.
-More importantly, a smaller number of iterations also suggests a better score function, since detection of promising and wrong trajectories occurs earlier within the fixed time interval $[0,T_{\mathrm{max}}]$.
-Figure 7c represents the variance around the probability estimate for both methods.
-As is already the case with the other panels, both methods are similar for large transition probabilities: the difference in the values of $(\sigma^{2})^{S}$ is less than $5\%$ of $(\sigma^{2})_{S}^{S}$ only for transition probabilities larger than $0.5$.
-$(\sigma^{2})_{R}^{S}$ is only smaller than $(\sigma^{2})_{S}^{S}$ for transition probabilities $p\geq 0.08$.
+The fewer iterations of TAMS-$R$ are consistent with its larger number of computed time steps: to converge in fewer iterations, trajectories must be branched earlier (within $[0,T_{\mathrm{max}}]$) so their resimulation involves the computation of more time steps.
+More importantly, this also suggests a better score function, since detection of promising and wrong trajectories occurs earlier within the time interval $[0,T_{\mathrm{max}}]$.
+Figure 7c represents the variance around the probability estimates for both methods.
+As in the other panels, both methods are similar for large transition probabilities: the difference in the values of $(\sigma^{2})_{R}^{S}$ and $(\sigma^{2})_{S}^{S}$ is less than $5\%$ of $(\sigma^{2})_{S}^{S}$ only for transition probabilities larger than $0.5$.
+$(\sigma^{2})_{R}^{S}$ is only smaller than $(\sigma^{2})_{S}^{S}$ for transition probabilities $\alpha_{R}^{S}\geq 0.08$.
 Otherwise, TAMS-$S$ systematically yields a smaller variance, with a difference that can be quite large: $\alpha_{R}^{S}$ is more than $50\%$ larger than $\alpha_{S}^{S}$ for half of the parameter values.
@@
 As we move towards the left of the diagram, the difficulty of the transfer learning does not allow the RC to adapt to new dynamics, and its estimates are increasingly scattered.
-For even lower probabilities, TAMS-$R$ cannot even converge in many of the $30$ runs but gives a correct probability estimate in others, leading to a very large variance.
+For even lower probabilities, TAMS-$R$ cannot even converge in many of the $30$ runs but gives a correct probability estimate in others, yielding a very large variance.
 This also explains why, to the very left of the diagram, there are some cases where $(\sigma^{2})_{R}^{S}<(\sigma^{2})_{S}^{S}$: in this case, there are so many TAMS runs that did not converge and returned a probability of 00 that the prediction of TAMS-$R$ becomes quite consistent.
-Unlike the variances presented in Fig.
-3, the variances shown in Fig.
-7c are computed over $30$ runs of TAMS-$R$, each using an independently trained RC.
+Note that the variances presented in Fig.
+3 were computed by running TAMS $30$ times for each of the $30$ independently trained RCs whereas the variance plotted in Fig.
+7c is directly computed from $30$ runs of TAMS, without freezing the RCs.
 5 Interpretability of the RC
-An important advantage of using RCs is that they can be interpreted.
-Each column of $\mathbf{W}_{\mathrm{out}}$ (see Appendix C) corresponds to a combination of powers of $S_{n},S_{ts},D$ and $t$.
-We can thus read in this matrix the terms that contribute the most to the estimated committor.
 Figure 8: Each panel shows, for a fixed $f_{\sigma}$, the terms with largest magnitude in the matrix $W_{\mathrm{out}}$ of the RCs.
-From the $30$ RCs trained in section 3.2, we extract the $10$ terms of $W_{\mathrm{out}}$ with largest magnitude.
-Their magnitude is then normalized to that of the largest term.
-We then only keep the terms that are most significant for more than $20$ out of $30$ RCs.
-Each curve correspond to a value of the parameter $\overline{E_{a}}$ and they show the average magnitude of each term across the $30$ RCs and the errorbars show the corresponding standard deviations.
-First, we focus on $\Phi_{R}^{F}$.
-In the case of F-transitions, we have run TAMS $30$ times for $30$ different RCs at $3$ values of $f_{\sigma}$ (see Sect.
-3.2).
-For each of the RCs, we select the $10$ values of $W_{\mathrm{out}}$ that have the largest amplitude (in absolute value), and we only keep the terms that are common in more than $20$ RCs.
-Their average magnitude (across the $30$ RCs) is normalized so that the largest is equal to $1$ in absolute value.
-Figure 8 presents these magnitudes with their standard deviation across the $30$ RCs.
-For each value of $f_{\sigma}$, if the most significant terms are similar for several values of $\overline{E_{a}}$, they are plotted together on the corresponding panel.
-From Fig.
-8, we see that for all values of $f_{\sigma}$, the most important terms of all RCs are very consistent: few error bars on the magnitude of every term are visible.
-This means that all RCs can extract a meaningful structure from the data.
-Furthermore, for each $f_{\sigma}$, the structure of the matrix $W_{\mathrm{out}}$ also converges as the transition probabilities become very small, as shown by the overlap of the curves obtained for several values of $\overline{E_{a}}$.
-These curves also allow for an analytical reconstruction of a suitable score function.
-For $f_{\sigma}=0.25$ and $f_{\sigma}=0.5$, $3$ interesting terms are highlighted: $S_{n}S_{ts}D^{2}$ (with a magnitude of $-1$) and $S_{n}^{2}D^{2}$ and $S_{ts}^{2}D^{2}$ (with a magnitude of about $0.5$).
+From the $30$ RCs trained in Sect.3b, we extract the $10$ terms of $W_{\mathrm{out}}$ with largest magnitude.
+Their magnitude is then normalized to that of the largest term and we only keep the terms that are most significant for more than $20$ out of $30$ RCs.
+Each curve corresponds to a value of the parameter $\overline{E_{a}}$ (the legend gives the corresponding transition probability computed with TAMS-$R$) and they show the average magnitude of each term across the $30$ RCs and the errorbars are the corresponding standard deviations.
+An important advantage of RCs is that we can read which term of their output matrix $W_{\mathrm{out}}$ contributed most to the committor estimation.
+In the case of F-transitions, we consider the $30$ runs of TAMS-$R$ performed at $f_{\sigma}=0.5,0.25$ and $0.05$.
+For each of the $30$ RCs, we select the $10$ largest terms of $W_{\mathrm{out}}$ in absolute value and normalize them, so that the largest coefficient in each RC has a magnitude of $1$.
+Then, we only keep the terms that are common in more than $20$ RCs (the results are robust when varying this number between $15$ and $25$), to compare how the most significant terms are ranked by each RC.
+Figure 8 presents these relative magnitudes with their standard deviation across the $30$ RCs.
+Each curve corresponds to a value of $\overline{E_{a}}$ (or equivalently to a grid cell in Fig.
+2a).
+For a given value of $f_{\sigma}$, they were sampled in the following way: starting from the smallest value of $\overline{E_{a}}$ such that $\alpha_{R}^{F}\geq 10^{-9}$, we retained one value of $\overline{E_{a}}$ out of $10$ such that $\alpha_{R}^{F}\leq 0.5$.
+By imposing a certain interval on the transition probabilities, we make sure that all the presented curves roughly correspond to similar dynamics of the model despite the different values of $(\overline{E_{a}},f_{\sigma})$.
+Figure 8 shows that, for all values of $f_{\sigma}$, the most important terms of all RCs are consistent: few error bars on the magnitude of every term are visible.
+All RCs thus extract a meaningful structure from the data and they all agree on the ranking of the most significant terms.
+Furthermore, for each $f_{\sigma}$, the structure of the matrix $W_{\mathrm{out}}$ converges as the transition probabilities become very small, as shown by the overlap of the curves obtained for several values of $\overline{E_{a}}$.
+These curves also provide an analytical expression of the learned score function.
+For $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$, $3$ interesting terms are highlighted: $S_{n}S_{ts}D^{2}$ (with a magnitude of $-1$) and $S_{n}^{2}D^{2}$ and $S_{ts}^{2}D^{2}$ (with a magnitude of about $0.5$).
 This naturally leads to a score function $\Phi_{R}^{F}\propto(S_{n}-S_{ts})^{2}D^{2}/2$.
-In both cases, we can also add the terms $(S_{ts}-S_{n})S_{ts}t^{2}/2$ and $(S_{n}^{2}+S_{ts}^{2}/2-S_{n}S_{ts})Dt/2$.
-These terms are not exactly contained in the coefficient of $W_{\mathrm{out}}$ because the coefficients of this matrix are not exactly $1$, $0.5$, or $0.25$.
-These analytical expressions give, nonetheless, a first estimate of an analytical formula for $\Phi_{R}^{F}$, for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$:
+In both cases, we can also add the terms $(S_{ts}-S_{n})S_{ts}t^{2}/2$ and $(S_{n}^{2}+S_{ts}^{2}/2-S_{n}S_{ts})Dt/2$ although these terms are not exactly contained in $W_{\mathrm{out}}$ because the coefficients of this matrix are not exactly $1$, $0.5$, or $0.25$.
+We still obtain a first estimate of an analytical result for $\Phi_{R}^{F}$, for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$:
 $\Phi_{R}^{F}(S_{n},S_{ts},D,t)\propto$ $a_{1}\frac{D^{2}}{2}(S_{n}-S_{ts})^{2}$ (5)
@@
 $+$ $a_{3}\frac{Dt}{2}\left(S_{n}^{2}+\frac{S_{ts}^{2}}{2}-S_{n}S_{ts}\right)$
-for certain ${\cal O}(1)$ coefficients $a_{i},\ i=1,2,3$, which have appropriate dimensions to make the expression consistent.
-This expression does not have an ”equal” sign because it is rather a proportionality relation, due to the normalization of $W_{\mathrm{out}}$.
-$\Phi_{R}^{F}$ can be compared to the structure of $\Phi_{S}^{F}$, driven by the term $(S_{n}-S_{ts})D^{2}$.
-Here, this precise term cannot be retrieved because the nonlinear part of the feature vector contains only monomials of degree $4$ (see Appendix C).
-The original implementation of the RC (Gauthier et al., 2021), which we followed, imposes a fixed choice of degree and we found $4$ to work best (Jacques-Dumas et al., 2023).
-Although $\Phi_{S}^{F}$ cannot be exactly reproduced, the RCs retrieve a similar dependence on $D^{2}$ and $(S_{n}-S_{ts})$.
-Most importantly, $\Phi_{R}^{F}$ involves time, which is what we should expect from a score function, since the transition must occur before $T_{\mathrm{max}}$.
+for certain ${\cal O}(1)$ coefficients $a_{i},\ i=1,2,3$ with appropriate dimensions to make the expression consistent.
+Due to the normalization of $W_{\mathrm{out}}$, this expression is only a proportionality relation.
+Moreover, the used implementation of the RC by Gauthier et al.
+(2021) imposes a fixed choice of degree of the terms in the nonlinear part of the feature vector and we found $p=4$ to work best (Jacques-Dumas et al., 2023).
+Although $\Phi_{S}^{F}\propto q_{N}\propto(S_{n}-S_{ts})D^{2}$ cannot be reproduced (because $q_{N}$ is of degree $3$), we can compare this expression to that of $\Phi_{R}^{F}$.
+The RCs retrieve a similar dependence on $D^{2}$ and $(S_{n}-S_{ts})$.
+This salinity difference drives the density gradient between the boxes n and ts, which, in turn, controls in this model the AMOC strength $q_{N}$ (Cimatoribus et al., 2014), implying that the salt advection drives the AMOC strength and that the salt advection feedback is key in the AMOC collapse.
+The presence of $(S_{n}-S_{ts})$ in all terms of $\Phi_{R}^{F}$ shows that the RC is able to capture the importance of this mechanism.
+$\Phi_{R}^{F}$ also involves time, which was to be expected from the score function.
+Indeed, TAMS returns a transition probability conditioned on a time horizon $T_{\mathrm{max}}$ so the score function is ideally time-dependent.
 In particular, the third term of Eq.
-6 is very interesting, as it combines $D$, $t$, and a term that looks like $(S_{n}-S_{ts})^{2}$, which would not have been possible if the degree of the monomials in the feature vector had been smaller.
-Finally, it is striking to notice that the expressions for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$ are very close.
-This is due in part to transfer learning, but, considering that the RCs are constantly independently retrained, the conservation of this structure across different values of $\overline{E_{a}}$ and $f_{\sigma}$ shows the robustness of this method.
-However, the RC is still able to adapt, as shown in the third panel of Fig.
-8 by the very different structure of $\Phi_{R}^{F}$ for $f_{\sigma}=0.05$.
+5 combines $D$, $t$, and a term looking like $(S_{n}-S_{ts})^{2}$, which would not be possible if the degree of the monomials was $p<4$.
+Finally, the expressions for $f_{\sigma}=0.5$ and $f_{\sigma}=0.25$ are strikingly close.
+This is partly due to transfer learning, but, considering that the RCs are constantly independently updated, the conservation of this structure across different values of $\overline{E_{a}}$ and $f_{\sigma}$ shows the robustness of this method.
+However, the different structure of $\Phi_{R}^{F}$ for $f_{\sigma}=0.05$ (third panel of Fig.
+8) shows that the RCs are still able to adapt.
 In this case, two monomials stand out: $S_{n}^{2}D^{2}$ and $-S_{ts}^{2}D^{2}$.
@@
 for certain ${\cal O}(1)$ coefficients $b_{i},\ i=1,2,3$.
-The estimated score function for $f_{\sigma}=0.05$ is different from the score obtained for much larger values of the noise amplitude, since a different noise induces different dynamics in the model.
-Thus, a data-driven score function is expected to depend on the model parameters.
-If we were able to solve the backward Kolmogorov equation (see Appendix B), the exact committor would depend on $f_{\sigma}$, which would make the structure of the score function evolve with $f_{\sigma}$.
-However, here such an analytical dependence on $f_{\sigma}$ cannot be retrieved and only its effect can be observed.
+The function $\Phi_{R}^{F}$ for $f_{\sigma}=0.05$ is different from the score function obtained for larger values of $f_{\sigma}$, which is not surprising as a data-driven score function is expected to depend on the model parameters.
+Moreover, $f_{\sigma}$ already appears in the equations of the committor (Lucente et al., 2022), so this function is expected to depend on it.
 Finally, despite the different structure of $\Phi_{R}^{F}$ for $f_{\sigma}=0.05$, its dependence on $(S_{n}-S_{ts})$ is conserved, although always multiplied by $(S_{n}+S_{ts})$, stressing more the importance of the salt-advection feedback.
@@
 Such an analysis may be very useful in more complex cases and larger-dimensional models to understand the possible structure of a good score function.
-We test this point by interpreting the output of $W_{\mathrm{out}}$ when trained to study the transition from the AMOC steady on state to the AMOC steady off-state (studied in Sect.
+We test this point by interpreting the output of $W_{\mathrm{out}}$ in the case of the AMOC S-transitions (see Sect.
 4).
@@
 8, obtained from the matrix $W_{\mathrm{out}}$ of $30$ RCs trained on an S-transition of the AMOC, for two values of $f_{\sigma}$: $0.5$ and $0.1$.
-For each value of $f_{\sigma}$, several curves are presented, corresponding to probability estimates $\alpha_{R}^{S}$ between $0.002$ and $0.1$.
+Each curve corresponds to a grid cell in Fig.
+6.
+For each value of $f_{\sigma}$, we display all curves for which the probability estimates $\alpha_{R}^{S}$ lies between $0.002$ and $0.1$.
+This corresponds to only two curves in panel b), because the jumps in transition probabilities between consecutive samples of $\overline{E_{a}}$ is larger for $f_{\sigma}=0.1$ than for $f_{\sigma}=0.5$ (as can be observed in Fig.
+6a).
 For both values of $f_{\sigma}$, all curves are very consistent, showing the ability of the RCs to capture some structure of the phase space and to converge, although they are constantly independently retrained.
-However, in both panels, the curves corresponding to the lowest transition probabilities have large error bars, showing the limits of transfer learning with a large difference between consecutive values of $\overline{E_{a}}$.
-In this case, the analytical formulas for the committor that can be extracted from these curves are:
+For $f_{\sigma}=0.5$, the curve corresponding to the largest $\alpha_{R}^{S}$ deviates from the others (on the right of panel a)), although it has large error bars.
+In both panels, the curves corresponding to the lowest transition probabilities also have large error bars, showing the limits of transfer learning with a large difference between consecutive samples of $\overline{E_{a}}$.
+In this case, the analytical results for the committor that can be extracted from these curves are:
 $f_{\sigma}=0.5:\ \Phi_{R}^{S}\propto$ $-c_{1}t^{2}\left(\frac{3S_{n}^{2}}{4}-S_{ts}^{2}+\frac{S_{n}S_{ts}}{2}\right)$ (7)
@@
 -- $d_{3}S_{ts}D^{2}t$
-for again specific coefficients $c_{i}$ and $d_{i}$.
-These expressions are difficult to relate to $\Phi_{S}^{S}$ because the latter relied purely on Euclidean distances in phase space rather than on some specific physical variables.
+for again specific ${\cal O}(1)$ coefficients $c_{i}$ and $d_{i}$.
+These expressions are difficult to relate to $\Phi_{S}^{S}$ because the latter only relies on Euclidean distances in phase space rather than specific physical variables.
 6 Discussion
@@
 On the one hand, Lucente et al.
-(2022) had already coupled Adaptive Multilevel Splitting (an algorithm almost similar to TAMS) to another machine learning technique, the Analogue Markov Chain, but the committor had to be estimated beforehand, thus creating an additional computational cost.
-On the other hand, direct committor function computation using machine learning has already been performed on complex climate data (Miloshevich et al., 2023b), but only with the objective of short-term probabilistic forecast.
-We have coupled TAMS to RC to show that the committor function could be learned at no additional cost than what is necessary to compute transition probabilities.
-We compared our results to a combination of TAMS experiments with a physics-informed score function and showed that the transition probabilities, the mean first-passage times and the transition paths were consistent.
-In the case of the F-transition, TAMS-$R$ also yields on average a similar variance in the results as TAMS-$S$, except for very low noise.
-Regarding the S-transition, both methods remain consistent for transition probabilities within $1000$ years down to $10^{-6}$.
-Furthermore, TAMS-$R$ consistently requires much fewer time steps to converge than TAMS-$S$, suggesting that in this case RCs can learn a more efficient score function than $\Phi_{S}^{S}$.
-Finally, we showed that analytical formulas can be derived for the learned committor are consistent in the case of the F-transition with the physics-informed score function.
-The next step is to study AMOC tipping in models of much higher dimensions, such as Earth Models of Intermediate Complexity (EMICs) or even Global Climate Models (GCMs).
-Another rare-event algorithm (Moral and Garnier, 2005; Tailleur and Kurchan, 2006; Lecomte and Tailleur, 2007) has already been applied to EMICs (Ragone et al., 2018; Ragone and Bouchet, 2021; Wouters et al., 2023) and to AMOC tipping in particular (Cini et al., 2023).
-It relies on other selection/deletion rules than a score function, but makes it more difficult to compute transition probabilities.
-Another interesting data-driven approach of AMOC tipping is the application of the edge-tracking algorithm by Mehling et al.
-(2024).
-It aims at finding the critical state through which an AMOC tipping would occur, using a limited number of simulations, and would thus be tractable in higher dimension.
-This approach provides insight into the dynamics of this system but not on quantities such as the probability of tipping or transition times.
+(2022) had already coupled Adaptive Multilevel Splitting (an algorithm similar to TAMS) to another machine learning technique, the Analogue Markov Chain, which estimated the committor beforehand, thus creating an additional computational cost.
+On the other hand, direct committor function estimation using machine learning has already been performed on complex climate data (Miloshevich et al., 2023b), but only for short-term probabilistic forecast.
+We showed that the committor function can be learned at no additional cost than needed to compute transition probabilities.
+We compared our results to a combination of TAMS with a physics-informed score function and showed that transition probabilities and mean first-passage times were consistent.
+In the case of F-transitions, TAMS-$R$ also yields on average a similar variance in the results as TAMS-$S$, except for very low noise.
+Regarding S-transitions, both methods remain consistent for transition probabilities within $1000$ years down to $10^{-6}$.
+Furthermore, TAMS-$R$ consistently converges in fewer iterations than TAMS-$S$, suggesting that, in this case, RCs can learn a more efficient score function than $\Phi_{S}^{S}$.
+Finally, we showed that the analytical interpretation of the learned committor is consistent in the case of the F-transition with the physics-informed score function.
+Concerning the AMOC, we confirmed that the probability of a collapse within a time frame of $100$ years for the F-transition and $1000$ years for the S-transition is not negligible for a large range of model parameters.
+Moreover, our method yields transition paths that are consistent with those obtained with a physics-informed score function.
+Finally, some of the terms of the estimated expression of the committor indicate that the RC captures the salt-advection feedback as key in the collapse process.
+The next step is to study AMOC tipping in higher-dimensional models like Earth Models of Intermediate Complexity (EMICs) or even Global Climate Models (GCMs).
+Another rare-event algorithm (Moral and Garnier, 2005; Tailleur and Kurchan, 2006; Lecomte and Tailleur, 2007) has already been applied to EMICs (Ragone et al., 2018; Ragone and Bouchet, 2021; Wouters et al., 2023) and AMOC tipping (Cini et al., 2024).
+However, it focuses on the sampling of rare events rather than computing transition probabilities.
+Another interesting data-driven approach is the application of the edge-tracking algorithm to AMOC tipping (Mehling et al., 2024).
+It finds the critical state through which an AMOC tipping would occur, using a limited number of simulations, and is thus tractable in higher dimensions.
+This approach provides insight into the dynamics of the system but does not quantify the tipping probability.
 Extending our method to complex climate models raises the question of its scaling with the dimension of the model.
-For TAMS-$R$, this question is twofold: the feasibility of simulating enough trajectories on the one hand, and the scaling of RC itself, on the other.
-Regarding the first point, TAMS is not more computationally demanding than any other rare-event algorithm so their recent successes in EMICs (Cini et al., 2023) are promising for TAMS.
+For TAMS-$R$, this question is twofold: the feasibility of simulating enough trajectories and the scaling of RC itself.
+Regarding the first point, TAMS is not more computationally demanding than any other rare-event algorithm so their recent successes in EMICs (Cini et al., 2024) are promising for TAMS.
 When it comes to GCMs (van Westen and Dijkstra, 2023), running TAMS does not seem feasible yet.
-However, the current implementation of the RC method poses a real scaling problem.
-Using monomials as a family of basis functions results in a feature vector whose size grows exponentially with the number of variables.
-We can only consider a very limited number of variables, but it raises in turn the question whether such simplification is sufficient to describe the dynamics of AMOC tipping in a GCM.
+However, the current implementation of the RC method poses a scaling problem.
+Due to the monomials used as basis functions, the size of the feature vector grows exponentially with the number of variables.
+We can only consider a very limited number of variables, which may be too simple to describe AMOC tipping in a GCM.
 Miloshevich et al.
-(2023a) recently showed that Convolutional Neural Networks (CNN) are suitable to learn the committor from maps of climate variables by reducing them into a lower-dimension vector.
-However, the problem remains that CNNs require a large amount of data to be trained.
+(2023a) showed that Convolutional Neural Networks (CNN) are suitable for learning the committor from maps of climate variables.
+However, training CNNs still requires a large amount of data.
 More generally, this raises the problem of the choice of basis functions for the RC.
 The same issue arises for another committor estimation method, the Dynamical Galerkin Approximation (Thiede et al., 2019; Finkel et al., 2021; Jacques-Dumas et al., 2023).
-It is interesting to note that Strahan et al.
+Strahan et al.
 (2023) recently used a Feedforward Neural Network to solve the Feynman-Kac formula for the committor.
-This approach only requires short trajectories such as those that can be obtained with TAMS but the issue of the amount of data and the training of the neural network remains.
+It only requires short trajectories such as those obtained with TAMS but may not solve the issue of the amount of data needed for training.
 The interpretability of the RC could also prove useful on the question of dimensionality.
-For example, we showed in Section 5 that $\mathbf{W}_{\mathrm{out}}$ contains forty basis terms but only about ten play a significant role in the final score function.
-If these can be detected early enough in the learning process, dropping them would greatly speed up the computation.
-More generally, neural networks are often considered black boxes and difficult to interpret.
-Here, it is possible, by estimating analytical formulas, to highlight important variables for the system dynamics and their relationship.
-It must be used with caution, but can serve as a useful starting point to study the dynamics of the system or to focus observations or experiments.
-The efficiency of TAMS is considered to rely on sufficient insight into the model’s dynamics (Lestang et al., 2018), but we show here that an efficient TAMS can also be used to obtain such useful insight.
-Finally, the coupling between TAMS and RC must be refined, especially the initialization and training process of the latter.
+We showed in Sect.
+5 that in the feature vector of the RC, only a dozen of the terms play a significant role in the estimate.
+If these can be detected early enough in the learning process, dropping the others would speed up the computation.
+More generally, neural networks are often considered to be black boxes and difficult to interpret.
+Here, important variables and their relationship can be highlighted, which may serve as a useful starting point to focus observations or experiments.
+The efficiency of TAMS may rely on sufficient insight into the model’s dynamics (Lestang et al., 2018), but we show here that an efficient TAMS can also provide such insight.
+Finally, the initialization and training process of the RC must be refined.
 In complex models, transition probabilities may only be computed for a single model setup, which makes transfer learning unusable.
 The RC could then be initialized from a linear score function that makes it easy to generate surrogate data.
-Since we might not be able to make a better guess of a prescribed score function, this could be a suitable starting point for the learning process.
 Another important issue is the memory of the RC.
-The current transfer learning process probably may retain too much information corresponding to very different system parameters, since we did not ”clean” its memory during the computations.
-This problem does not only arise with transfer learning, and we would also encounter it if the initialization of the RC is very far from the actual committor.
-A possible solution could be to reinitialize the RC at certain milestones, once a certain number of transitions have been reached, thus using ”better” data.
+The current transfer learning process may retain too much information corresponding to very different system parameters since we did not ”clean” its memory during the computations.
+The same issue may arise if the initialization of the RC is very far from the actual committor.
+A possible solution is to reinitialize the RC at certain milestones, once a certain number of transitions have been reached, thus using ”better” data.
 Acknowledgements.
@@
   2. 2.
-    Generate $N$ trajectories $\left(\mathbf{X}^{(1)},\dots,\mathbf{X}^{(N)}\right)$ starting in set $A$ until $T_{\mathrm{max}}$ or until they reach set $B$.
-Repeat steps $3-9$ until $N-n_{c}$ trajectories have reached $B$ (or $N$ if $n_{c}=1$).
+    Simulate $N$ trajectories $\left(\mathbf{X}^{(1)},\dots,\mathbf{X}^{(N)}\right)$ starting in set $A$ until $T_{\mathrm{max}}$ or until they reach set $B$.
+Repeat steps $3-8$ until at least $N-n_{c}$ (or $N$ if $n_{c}=1$) trajectories have reached $B$.
   3. 3.
@@
     For trajectory $i$, the maximum value of the score function is called $Q_{i}$.
-  5. 5.
 Let $I_{k}=\{i\in[1,N]\ |\ Q_{i}\in\{Q^{*}_{1},...,Q^{*}_{n_{c}}\}\}$, where $\{Q^{*}_{j}\}_{j\in[1,n_{c}]}$ are the $n_{c}$ lowest values of $Q_{i}$.
 Since the system is discrete, $I$ may contain more than $n_{c}$ elements.
+  5. 5.
+    Set $k=k+1$ then $w_{k}=\left(1-\frac{\#I_{k}}{N}\right)w_{k-1}$. For every $i\in I_{k}$, repeat steps $7-9$.
   6. 6.
-    Set $k=k+1$ then $w_{k}=\left(1-\frac{\#I_{k}}{N}\right)w_{k-1}$. For every $i\in I_{k}$, repeat steps $7-9$.
+    Randomly select a trajectory $\mathbf{X}^{(r)}$ (such that $r\notin I_{k}$). We call $\tau$ the first time so that $\Phi(\mathbf{X}^{(r)}(\tau))\geq Q_{i}$.
   7. 7.
-    Randomly select one of the remaining trajectories, say $\mathbf{X}^{(r)}$ (such that $r\notin I_{k}$). We call $\tau$ the first time so that $\Phi(\mathbf{X}^{(r)}(\tau))=Q_{i}$.
+    Set $\mathbf{X}^{(i)}([1,\tau])=\mathbf{X}^{(r)}([1,\tau])$.
   8. 8.
-    Set $\mathbf{X}^{(i)}([1,\tau])=\mathbf{X}^{(r)}([1,\tau])$.
-  9. 9.
     Simulate the rest of $\mathbf{X}^{(i)}$ starting from $\mathbf{X}^{(r)}(\tau)$ until time $T_{\mathrm{max}}-\tau$ or until reaching $B$.
-We call $N_{B}$ the final number of trajectories that contain a transition, and the total number of iterations is called $K$.
-The estimated transition probability from $A$ to $B$ before $T_{\mathrm{max}}$ is given by the equation (Rolland and Simonnet, 2015; Rolland, 2022):
-$\hat{q}=\frac{N_{B}}{N}w_{K}=\frac{N_{B}}{N}\prod_{k=1}^{K}w_{k}=\frac{N_{B}}{N}\prod_{k=0}^{K-1}\left(1-\frac{\#I_{k}}{N}\right)$ (9)
-The optimal variance, obtained when the score function is the committor function (see Appendix B), reads:
-$\sigma^{2}_{\mathrm{id}}=\frac{<\hat{q}>^{2}}{N}\left(<K>\frac{n_{c}}{N-n_{c}}+\frac{N-<N_{B}>}{<N_{B}>}\right)$ (10)
-where <>$\cdot$>< ⋅ > represents an average over the $M$ runs of TAMS.
+Let $N_{B}$ the number of trajectories that reached $B$ and $K$ and the number of iterations.
+The estimated transition probability from $A$ to $B$ before $T_{\mathrm{max}}$ is given by (Rolland and Simonnet, 2015; Rolland, 2022):
+$\hat{\alpha}=\frac{N_{B}}{N}w_{K}=\frac{N_{B}}{N}\prod_{k=1}^{K}w_{k}=\frac{N_{B}}{N}\prod_{k=0}^{K-1}\left(1-\frac{\#I_{k}}{N}\right)$ (8)
+The optimal variance, obtained when the score function is the committor function reads:
+$\sigma^{2}_{\mathrm{id}}=\frac{<\hat{\alpha}>^{2}}{N}\left(<K>\frac{n_{c}}{N-n_{c}}+\frac{N-<N_{B}>}{<N_{B}>}\right)$ (9)
+where <>$\cdot$>< ⋅ > represents an average over $M$ runs of TAMS.
 […]
-Consider a stochastic dynamical system of dimension $d$ and two subsets $\mathcal{E}$ and $\mathcal{F}$ of its phase space.
-If the drift of the system is called $a$ and its diffusivity $D$, then the adjoint $\mathcal{L}$ of the Fokker-Planck operator is the infinitesimal generator of the stochastic process and defined as:
-$\mathcal{L}=\frac{\partial}{\partial t}(\cdot)-\sum_{i=1}^{n}a_{i}(\mathbf{x})\frac{\partial}{\partial\mathbf{x}_{i}}(\cdot)-\sum_{i,j=1}^{n}D_{ij}(\mathbf{x})\frac{\partial}{\partial\mathbf{x}_{i}\partial\mathbf{x}_{j}}(\cdot),$ (11)
-Formally, the committor function, as solution to the backward Kolmogorov equation, solves the Dirichlet problem (Lucente et al., 2022):
-$\mathcal{L}\Phi(t,\mathbf{x})$ $=0\ \forall(t,\mathbf{x})\in\mathbb{R}\times(\mathcal{E}\cup\mathcal{F})^{c}$ (12)
-$\Phi(t,\mathbf{x})$ $=1\ \forall(t,\mathbf{x})\in\mathbb{R}\times\mathcal{E}$
-$\Phi(t,\mathbf{x})$ $=0\ \forall(t,\mathbf{x})\in\mathbb{R}\times\mathcal{F}$
-The committor function $\Phi(t,\mathbf{x})$, as defined by the system (12) and its boundary conditions, also corresponds to the probability that a trajectory $\mathbf{X}(t)$ starting on $\mathbf{x}$ reaches the set $\mathcal{E}$ before $\mathcal{F}$.
-More precisely, for any set $\mathcal{G}\in\mathbb{R}^{n}$, the first passage time in $\mathcal{G}$ is defined as:
-$\tau_{\mathcal{G}}=\min\{t\ |\ \mathbf{x}(t)\in\mathcal{G}\}$ (13)
-The committor is defined as the probability that the first-passage time in $\mathcal{E}$ is lower than that in $\mathcal{F}$:
-$\Phi(t,\mathbf{x})=\mathbb{P}(\tau_{\mathcal{E}}<\tau_{\mathcal{F}}\ |\ \mathbf{X}(0)=\mathbf{x})$ (14)
-[…]
-Reservoir Computing (Jaeger, 2001; Lukoševičius and Jaeger, 2009) is a kind of neural network that is designed to better represent dynamical systems and temporal processes than classical feedforward neural networks, without using gradient descent for training.
-Reservoir Computing aims at transforming a nonlinear problem in the original phase space into a linear problem by linearly embedding input data into a larger-dimensional space.
-Then a linear regression is performed in this feature space.
-The main idea is that the matrix representing this space is random and fixed; only the linear regression parameters are learned during training.
-Although this neural network acts as a universal approximator (Gonon and Ortega, 2020), its use encounters certain problems, such as the randomness of the reservoir: different realizations of a random network can greatly change the performance of the reservoir computer.
-To solve this issue, Gauthier et al.
-(2021) developed another universal approximator (Gonon and Ortega, 2020; Hart et al., 2021) that they call “Next-Generation” Reservoir Computing (here called RC).
-The difference is that the input data is now embedded into a larger-dimensional space using a family of nonlinear basis functions instead of a random matrix.
+Reservoir Computing (Jaeger, 2001; Lukoševičius and Jaeger, 2009) is a kind of neural network designed to represent dynamical systems and temporal processes better than feedforward networks.
+The idea is to transform a nonlinear problem in the original phase space into a linear problem by embedding input data into a larger-dimensional space where linear regression is performed.
+Classical reservoir computers are universal approximators (Gonon and Ortega, 2020) but pose certain problems.
+For instance, their performance can vary depending on the realizations of the reservoir, which is usually a fixed random matrix.
+The “Next-Generation” Reservoir Computing (here abbreviated RC) developed by Gauthier et al.
+(2021) is also an universal approximator (Gonon and Ortega, 2020; Hart et al., 2021) but embeds input data into a large-dimensional space using a family of nonlinear basis functions instead of a random matrix.
 We provide below only the important steps to build an RC for our problem.
@@
 (2023) for the setup and choice of parameters.
-At each time step, the RC is given a vector $\mathbf{x}\in\mathbb{R}^{d}$ that contains here only $S_{n}$, $S_{ts}$, $D$ and time $t$.
-The family of non-linear functions consists of products of power $p$ of these variables.
-The resulting feature vector $\mathbf{\tilde{x}}\in\mathbb{R}^{M}$, where $M=40$, is much smaller than in classical reservoir computing (where its dimension would be approximately $1,000$), making the training of RC less computationally intensive.
-Finally, the output matrix $\mathbf{W}_{\mathrm{out}}\in\mathbb{R}^{2\times M}$ maps $\mathbf{\tilde{x}}$ to the target output $\mathbf{y}\in\mathbb{R}^{2}$.
-Since we do not have access to any values of the committor, these cannot be used as target output for the linear regression.
-Instead, the only accessible information is whether $\mathbf{x}$ first leads the trajectory to $A$ or to $B$ or if $T_{\mathrm{max}}$ is reached before any of these sets.
-So the target output is $\mathbf{y}=(y_{1},y_{2})$ where $y_{1}=1$ if $A$ or $T_{\mathrm{max}}$ is reached first and 00 otherwise ; $y_{2}=1$ if $B$ is reached first and 00 otherwise.
-A Softmax operation then transforms these predictions finally into probabilities, namely: (”probability to reach $A$ or $T_{\mathrm{max}}$ first”, ”probability to reach $B$ first”).
-The latter corresponds to the committor function.
-Because it is a simple linear regression that transforms $\tilde{\mathbf{x}}$ into $\mathbf{y}$, an analytical expression of $\mathbf{W}_{\mathrm{out}}$ can be derived (Lukoševičius and Jaeger, 2009; Gauthier et al., 2021), only depending on the feature vectors $\mathbf{\tilde{X}}$ and associated labels $\mathbf{Y}$:
+The RC receives at each time step a vector $\mathbf{x}\in\mathbb{R}^{4}$ containing $S_{n}$, $S_{ts}$, $D$ and the time $t$.
+The family of non-linear functions consists of all products of power $p=4$ of these variables.
+The resulting feature vector $\mathbf{\tilde{x}}\in\mathbb{R}^{d}$, with $d=40$ is much smaller than in classical reservoir computing (where its dimension would be approximately $1,000$), making the training of RC more efficient.
+The output matrix $\mathbf{W}_{\mathrm{out}}\in\mathbb{R}^{2\times d}$ maps $\mathbf{\tilde{x}}$ to the target output $\mathbf{y}\in\mathbb{R}^{2}$.
+Since the committor is inaccessible, its values cannot be used as target for the linear regression.
+We only know whether $\mathbf{x}$ first leads the trajectory to $A$ or to $B$ or if $T_{\mathrm{max}}$ is reached before any of these sets.
+The target output is $\mathbf{y}=(y_{1},y_{2})$ where $y_{1}=1$ if $A$ or $T_{\mathrm{max}}$ is reached first and 00 otherwise ; $y_{2}=1$ if $B$ is reached first and 00 otherwise.
+A Softmax operation then transforms these predictions into probabilities, namely: (”probability to reach $A$ or $T_{\mathrm{max}}$ first”, ”probability to reach $B$ first”).
+The second is an estimate of the committor.
+Since $\tilde{\mathbf{x}}$ and $\mathbf{y}$ are matched through linear regression, an analytical expression of $\mathbf{W}_{\mathrm{out}}$ can be derived (Lukoševičius and Jaeger, 2009; Gauthier et al., 2021), only depending on the feature vectors $\mathbf{\tilde{X}}$ and associated labels $\mathbf{Y}$:
 $\mathbf{W}_{\mathrm{out}}=\mathbf{Y}\mathbf{\tilde{X}}^{\mathrm{T}}(\mathbf{\tilde{X}}\mathbf{\tilde{X}}^{\mathrm{T}}+\gamma\mathbf{I})^{-1}$ (10)
-where $\mathbf{I}$ is the identity matrix and $\gamma$ is the Tikhonov regularization parameter. In our setup, $\mathbf{\tilde{X}}$ depends only on $\gamma=10^{-4}$ and on $p=4$.
-Finally, an important advantage of using RC is their ability to be trained online.
-Here, we would like to use as little data as possible to train the RC, so we use only the data provided by TAMS and retrain the RC at every iteration of the algorithm.
-Let $\mathbf{\tilde{X}}_{k}$ and $\mathbf{Y}_{k}$ be all the feature vectors and all the labels we used up to iteration $k$.
-If new data become available, we can add the newly computed feature vectors and labels (say $\mathbf{\tilde{x}}$ and $\mathbf{y}$) as new columns of $\mathbf{\tilde{X}}_{k}$ and $\mathbf{Y}_{k}$.
-The matrices $\mathbf{Y}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}$ and $\mathbf{\tilde{X}}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}$ can then be easily updated:
+where $\mathbf{I}$ is the identity matrix and $\gamma$ is the Tikhonov regularization parameter.
+Finally, an important advantage of using RC is their ability to be trained online, for instance to update $W_{\mathrm{out}}$ at each iteration of TAMS.
+Let $\mathbf{\tilde{X}}_{k}$ and $\mathbf{Y}_{k}$ be all the feature vectors and all the labels used up to iteration $k$.
+If new data become available, the newly computed $\mathbf{\tilde{x}}$ and $\mathbf{y}$ become new columns of $\mathbf{\tilde{X}}_{k}$ and $\mathbf{Y}_{k}$.
+The matrices $\mathbf{Y}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}$ and $\mathbf{\tilde{X}}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}$ are then updated:
 $\mathbf{Y}_{k+1}\mathbf{\tilde{X}}_{k+1}^{\mathrm{T}}$ $=\mathbf{Y}_{k}\mathbf{\tilde{X}}_{k}^{\mathrm{T}}+\mathbf{y}\mathbf{\tilde{x}}^{\mathrm{T}}$ (11)